### 18 Mar 2026 v2

# Forecasting and Manipulating the Forecasts of Others

Abstract
In strategic environments with private information, evaluating a change in policy requires predicting how the equilibrium responds—but when actions reshape opponents’ signals, each agent’s optimal response depends on an infinite hierarchy of beliefs about beliefs [28] that has resisted exact analysis for four decades. We provide the first exact equilibrium characterization of finite-player continuous-time linear quadratic games with endogenous signals. Conditioning on primitive Brownian shocks rather than the physical state—a dynamic analogue of Harsanyi’s common-prior construction—collapses the belief hierarchy onto deterministic impulse response maps, reducing Nash equilibrium to a deterministic fixed point with no truncation and no large-population limit. The characterization yields an explicit information wedge that prices the marginal value of shifting opponents’ posteriors. The wedge vanishes precisely when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters, and provides a closed-form mapping from information primitives to equilibrium outcomes.
1 Introduction
In strategic environments with private information, evaluating a change in policy—a disclosure mandate, a shift in the precision of public communication, a redesign of market transparency—requires understanding how the equilibrium responds, not merely how agents behave under existing rules. When a regulator alters the information structure, each agent’s optimal action depends on their beliefs about the state and on their beliefs about how others are filtering, and both change with the policy. The filtering environment and the strategic equilibrium are jointly determined.
This joint determination is the central obstacle. Consider a central bank shifting from quarterly to real-time disclosure of its output gap estimates. Each firm’s pricing decision depends on its belief about the gap and on its belief about how other firms filter the bank’s announcements. Under the new policy, signal precision changes; firms filter differently, revise their forecasts of one another’s pricing, and equilibrium inflation dynamics shift in tandem. Existing tools cannot trace this chain because the Townsend hierarchy makes the mapping from signal precision to equilibrium outcomes intractable. The characterization in this paper makes it explicit.
The same obstruction arises across a wide range of fields. In market microstructure, prices aggregate dispersed information and trading itself is informative. Dynamic information design faces the same difficulty: the mapping from signal structures to equilibrium outcomes becomes intractable once filtering and equilibrium interact. Contract theory encounters it when multiple agents exert unobservable effort on a shared outcome and draw inferences from it. So does decentralized control of autonomous vehicle fleets and multi-robot systems with misaligned operators, wherever agents observe a shared state through private channels and actions reshape what others sense. Existing approaches to the decentralized control problem typically assume cooperative teams or a common information base; the framework here requires neither.
Townsend [28] crystallized the difficulty: in a game with private signals, each agent must forecast others’ forecasts of the state, generating an infinite hierarchy of conditional expectations. Sargent [26] showed that even in linear–Gaussian environments the hierarchy prevents finite-dimensional recursive analysis, because each agent’s action enters the state dynamics, entangling every agent’s past actions, beliefs, and signals in a way the Kalman filter cannot close.
The conceptual move parallels Harsanyi’s [10] resolution of the belief hierarchy in incomplete-information games. Harsanyi replaced an infinite regress of beliefs about beliefs with a common prior over a type space from which all higher-order beliefs follow by conditioning. We replace the Townsend regress with the common probability space of primitive shocks $W=(W^{0},\ldots,W^{n})$, from which all higher-order conditional expectations follow by linear projection onto deterministic impulse response maps with no truncation. The analogy is imperfect in a revealing way: Harsanyi’s types are drawn once and the information partition is fixed, whereas here noise unfolds continuously and actions reshape signals, so the closure requires not just a common prior but a fixed point—the maps governing how projections evolve must be consistent with the strategy profile they induce (Figure 1).
In the single-agent problem the conditional mean of the state is a sufficient statistic, but the state estimate is a linear functional of the noise estimate, not the reverse. In the multi-agent setting the state estimate fails to close the equilibrium because it entangles exogenous shocks with endogenous responses—precisely the Townsend hierarchy. Estimating the noise path avoids this entanglement. We call the path $\widehat{W}_{t}^{i}(\cdot):=\mathbb{E}[W_{\cdot}\mid\mathcal{F}_{t}^{i}]$ the player’s noise-state; optimal controls are affine functionals of it, just as single-agent controls are affine in the state estimate.
The main results establish that the best response amongst all admissible actions to any profile of linear strategies is itself linear. The equilibrium characterization produces an explicit information wedge $\mathcal{V}^{i}_{t}$, which prices the marginal value to player $i$ of shifting opponents’ posteriors. It vanishes identically when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters. The propagation $\Delta P^{i}\to\Delta\widetilde{X}^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta D^{i}$ provides an explicit mapping from information primitives to equilibrium outcomes that a model with fixed filtering cannot deliver: because information enters through precision paths and affects the equilibrium only through deterministic impulse response maps, a principal can trace how disclosure precision propagates into equilibrium payoffs without simulating higher-order beliefs.
(i) Candidate Opponents use Noise-State Linear controls (ii) Closure State & filters have det. response maps (iii) Deviations Fixed maps $\Rightarrow$ det. propagation (iv) Best response FOC $\Rightarrow$ control noise-state linear (v) Fixed point BR map closes $\Longleftrightarrow$ Nash eq.
Figure 1: The Noise-State Linear fixed-point loop. (i) Assume opponents use Noise-state Linear controls. (ii) The state and filtering equations have deterministic impulse responses. (iii) Under fixed strategy maps, unobserved deviations propagate deterministically. (iv) The maximum principle produces a best response that is itself noise-state linear, closing the loop. (v) A fixed point of this best-response map is a Nash equilibrium.
1.1 Related Literature
The joint determination of filtering and equilibrium has been approached from several directions, each achieving tractability by restricting or removing the informational externality that makes the problem hard.
Higher-order expectations and the social value of information.
The infinite regress of “forecasting the forecasts of others” was identified by Townsend [28] and shown by Sargent [26] to resist finite-dimensional recursive analysis even in linear–Gaussian environments. Morris and Shin [20] demonstrated that the social value of public information can be negative when agents overweight common signals—a static manifestation of the belief-manipulation channel that arises dynamically through the information wedge $\mathcal{V}^{i}_{t}$. Our Proposition 3.6 provides a dynamic counterpart: Section 4 shows that nearly all welfare gains from pooling come from eliminating the strategic channel rather than improving state estimation, consistent with the Morris–Shin intuition that the coordination externality, not statistical efficiency, drives welfare losses.
Frequency-domain methods and exogenous-signal models.
A substantial literature solves dispersed-information models in the frequency domain [9, 14, 24], culminating in Huo and Takayama [12], who obtain finite-state representations for equilibrium aggregates under exogenous ARMA signals. These methods succeed because the filtering environment is invariant to changes in control laws—the spectral factorization can be solved independently of the policy. When controls enter the state dynamics the transfer function depends on the policy, breaking this separation; the present paper addresses this complementary case.
Common-information approaches.
A powerful technique in decentralized control conditions on a common information base shared by all agents to obtain a recursive reduction [21]. The approach handles asymmetric information but resolves it through a coordinator’s beliefs; the channel by which player $i$’s action shifts player $k$’s posterior, which shifts $k$’s action, which shifts $i$’s signal, is not captured.
Team theory.
The decentralized decision problem with shared objectives originates with Marschak and Radner [19]; Radner [23] established certainty equivalence and linear optimal policies for linear quadratic gaussian (LQG) teams. Team-theoretic solutions bypass the belief hierarchy because a social planner with aligned objectives solves a single-agent control problem—the fixed-point structure of Section 3 never arises. The information wedge $\mathcal{V}^{i}_{t}$ is an equilibrium object: it prices the value of shifting an opponent’s posterior, which requires distinct objectives to be well-defined.
Mean-field LQG games.
Mean-field games take large-population limits in which no single agent’s action measurably affects any other agent’s signal [6, 11]. Carmona and Delarue [7] provide a comprehensive two-volume treatment including LQG specifications with common noise. The bilateral informational externality is absent by construction in the mean-field limit; $\mathcal{V}^{i}_{t}$ is a finite-$n$ object that vanishes in any limit where individual actions become negligible.
Truncation of the belief hierarchy.
Nimark [22] approximates the higher-order belief hierarchy by truncation at finite order—the closest existing method in spirit to the present paper. The deterministic impulse-response system derived here provides an exact benchmark against which truncation error can be assessed.
Applied continuous-time filtering games.
Kyle [15], Back [2], and Foster and Viswanathan [8] solve strategic trading problems with private information, each exploiting special structure—a single informed trader, correlated-signal symmetry—to close the belief hierarchy through model-specific conjectures. Sannikov [25] solves a continuous-time principal–agent problem by showing that the agent’s continuation value serves as a sufficient statistic for the principal’s problem, yielding a recursive characterization despite the agent’s private action. The technique relies on the bilateral structure: the principal’s belief about a single agent’s hidden action can be summarized by a one-dimensional state variable, and the agent has no beliefs about other agents to forecast. With multiple agents exerting unobservable effort on a shared outcome—the moral-hazard analogue of our setting—each agent draws inferences from the shared output about others’ effort, generating the same belief hierarchy that arises here. The noise-state framework extends naturally to this setting: each agent’s hidden effort enters the output dynamics exactly as controls enter the state in (2.1), and the information wedge prices the value of shifting the principal’s and co-agents’ posteriors about one’s effort. The present paper provides a general framework that nests these as special cases: in the Kyle–Back embedding (Section 5), the information wedge corresponds to the strategic component of price impact, and the characterization extends to multi-trader settings where guess-and-verify breaks down.
Information design.
Kamenica and Gentzkow [13] established Bayesian persuasion; Bergemann and Morris [3, 4] introduced Bayes correlated equilibrium and survey the field, identifying dynamic decentralized settings as a central open problem. Bergemann, Heumann, and Morris [5] provide precision comparative statics in static coordination games analogous to the propagation $\Delta P^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta J^{i}$ derived here. The difficulty in extending these results dynamically is that filtering and equilibrium interact; our characterization provides the equilibrium mapping that dynamic information design requires as an input.
Rational inattention.
Sims [27] introduced rational inattention; Maćkowiak and Wiederholt [17] showed that firms optimally attend to idiosyncratic over aggregate conditions, generating sluggish price adjustment (see Maćkowiak, Mat́ejka, and Wiederholt [18] for a survey). These models take the strategic environment as given when solving the attention problem; Corollary 3.7 and the FOC (C.9) extend the framework to settings where the marginal value of precision depends on the equilibrium through the information wedge.
2 The Decentralized LQG Game
Informal description of the game.
The class of games we study has three features: (i) multiple agents share a common evolving state but observe it through private noisy channels; (ii) each agent’s action feeds back into the state dynamics and thereby into every other agent’s signal process; and (iii) agents minimize individual quadratic objectives over a finite horizon. Because actions enter the state, they simultaneously serve the physical objective and reshape what opponents observe, making each agent’s problem jointly strategic and informational. Features (i)–(ii) generate the infinite belief hierarchy identified by Townsend; feature (iii) provides the linear–Gaussian structure that makes it tractable—the same structure that underlies the workhorse models of [28, 26, 20] and the mean-field LQG literature [7], and that serves as the exact benchmark against which approximations can be assessed.
2.1 Primitives
Consider $n$ players interacting over a finite time horizon $[0,T]$ on a filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$.
State dynamics.
The state $X_{t}\in\mathbb{R}^{d}$ evolves as
$dX_{t}=\bigl(A(t)X_{t}+\sum_{i=1}^{n}D_{t}^{i}\bigr)dt+\Sigma(t)\,dW_{t}^{0},\qquad X_{0}=x_{0},$ (2.1)
where $A(t)\in\mathbb{R}^{d\times d}$ is deterministic and bounded, $W^{0}$ is a $d$-dimensional Brownian motion, $\Sigma(t)\in\mathbb{R}^{d\times d}$ is deterministic and bounded, and $D_{t}^{i}\in\mathbb{R}^{d}$ is player $i$’s control.
Remark 2.1 (Control matrices and dimensions).
All results extend to $D_{t}^{i}=B_{i}(t)U_{t}^{i}$ with control $U^{i}_{t}$ by appropriately modifying the cost function and first order conditions.
Observations.
Player $i$ does not observe $X_{t}$ directly. Instead, they observe a noisy measurement process $Y^{i}_{t}\in\mathbb{R}^{d}$ given by
$dY^{i}_{t}=\sqrt{P^{i}(t)}X_{t}\,dt+dW^{i}_{t},\quad Y^{i}_{0}=0,$ (2.2)
where: [^1] The square-root parameterization is convenient for interpreting $P^{i}(t)$ as a precision matrix, but all results extend to observation channels of the form $dY^{i}_{t}=H^{i}(t)X_{t}\,dt+E^{i}(t)\,dW_{t}$ with arbitrary deterministic $H^{i}(t)\in\mathbb{R}^{d_{i}\times d}$ and $E^{i}(t)E^{i}(t)^{\top}=I$, including correlated observation noises across players and signals that weight different state coordinates asymmetrically. All kernel formulas remain verbatim under this generalization.
  • $W^{i}_{t}\in\mathbb{R}^{d}$ is a standard Brownian motion (player $i$’s observation noise),
  • $P^{i}(t)\in\mathbb{R}^{d\times d}$ is symmetric positive semidefinite, and $\sqrt{P^{i}(t)}$ denotes its unique symmetric positive semidefinite square root.
  • All Brownian motions $\{W^{0},W^{1},\ldots,W^{n}\}$ are mutually independent.
Notation for primitive shocks.
We collect all primitive noises into a single vector $W_{t}:=(W^{0}_{t},\dots,W^{n}_{t})^{\top}\in\mathbb{R}^{(n+1)d}$. For each $j\in\{0,\dots,n\}$, the block selector $E^{j}\in\mathbb{R}^{d\times(n+1)d}$ extracts the $j$-th channel: $E^{j}W_{t}=W^{j}_{t}$. The block projector $\Pi^{i}:=E^{i,\top}E^{i}$ is the identity on the $i$-th block and zero elsewhere: $\Pi^{i}v=(0,\dots,0,v^{i},0,\dots,0)$. Player $i$ directly observes $\Pi^{i}W$ and must infer $(I-\Pi^{i})W$ from drift-based learning.
Player $i$’s information at time $t$ is their observation history, $\mathcal{F}^{i}_{t}:=\sigma(Y^{i}_{s}:s\leq t)$. An admissible control is $\mathcal{F}^{i}_{t}$-adapted with $\mathbb{E}[\int_{0}^{T}\|D^{i}_{t}\|^{2}\,dt]<\infty$, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$.
Unresolved impulse response.
For each player $i$, define
$\widetilde{X}^{i}_{t}(u):=X_{t}(u)-\widehat{X}^{i}_{t}(u),$
the component of the state’s impulse response to the $u$-th shock that player $i$ has not yet resolved at time tt—large when the shock is recent or the observation channel weak, and shrinking as observations accumulate. Formally, $\widetilde{X}^{i}_{t}(u)=\operatorname{Cov}(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i})/du$; it serves as the filtering gain in the noise-state update (Theorem 3.1), with its algebraic representation in terms of $F^{i}$ given in (A.12).
Objectives.
The linear state-cost coefficient $G^{X,i}_{t}=\bar{G}^{X,i}_{t}+\int_{0}^{t}G^{X,i}_{t}(r)\,dW_{r}$ is a stochastic process with deterministic mean $\bar{G}^{X,i}_{t}\in\mathbb{R}^{d}$ and deterministic kernel $G^{X,i}_{t}(r)\in\mathbb{R}^{d\times(n+1)d}$; deterministic coefficients are the special case $G^{X,i}_{t}(r)\equiv 0$. Each player $i$ incurs the random cost
$\begin{split}\mathcal{C}^{i}(D^{i};D^{-i}):=&\int_{0}^{T}\left(X_{t}^{\top}G^{XX,i}(t)X_{t}+2(G^{X,i}_{t})^{\top}X_{t}+G^{i}(t)+(D^{i}_{t})^{\top}G^{DD,i}(t)D^{i}_{t}\right)dt\\ &+X_{T}^{\top}G^{XX,i}(T)X_{T}+2(G^{X,i}_{T})^{\top}X_{T},\end{split}$ (2.3)
where $G^{XX,i}(t)$ and $G^{DD,i}(t)$ are symmetric positive semidefinite matrices (with $G^{DD,i}(t)$ positive definite), and $D^{-i}:=(D^{j})_{j\neq i}$ denotes opponents’ controls. Each player $i$ seeks to minimize
$J^{i}(D^{i};D^{-i}):=\mathbb{E}\bigl[\mathcal{C}^{i}(D^{i};D^{-i})\bigr].$ (2.4)
Opponents’ controls enter player $i$’s cost only through $X_{t}$, which they move via (2.1) and which player $i$ observes through (2.2).
2.2 Nash Equilibrium
Because each player’s drift depends on opponents’ estimates $\widehat{X}_{t}^{k}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{k}]$, optimal actions generate an infinite hierarchy of conditional expectations. The following equilibrium concept fixes opponents’ strategy maps rather than their realized actions, which is essential for the closure results in Section 3.
Definition 2.1 (Nash equilibrium in private-signal strategies).
Throughout, a player’s choice is a strategy map: for each $t\in[0,T]$, the action $D_{t}^{i}$ is a (progressively measurable) functional of the private observation history $Y^{i}_{\cdot\wedge t}$. We write $D_{t}^{i}=D_{t}^{i}[Y^{i}]$, where the bracket notation $D_{t}^{i}[\cdot]$ denotes the map from the observation path $(Y^{i}_{s})_{s\leq t}$ to the action at time $t$. We identify a strategy with its induced adapted control process, assuming $\mathbb{E}\int_{0}^{T}\|D_{t}^{i}\|^{2}\,dt<\infty$.
A profile $D^{*}=(D^{*,1},\ldots,D^{*,n})$ is a Nash equilibrium if for every player $i$ and every admissible alternative strategy map $\widetilde{D}^{i}[\cdot]$,
$J^{i}(D^{*,i};D^{*,-i})\leq J^{i}(\widetilde{D}^{i};D^{*,-i}),$
where the inequality compares the costs induced by the corresponding strategy profiles. Equivalently: in a unilateral deviation by $i$, opponents’ maps are held fixed and are applied to the deviated signal paths.
In a unilateral deviation by $i$, opponents’ strategy maps $Y^{k}_{\cdot\wedge t}\mapsto D_{t}^{k}[Y^{k}]$ are held fixed and applied to the perturbed observation histories; in the noise-state linear class these maps are parameterized by deterministic maps that remain unchanged off equilibrium.
The requirement that opponents’ strategy maps be held fixed under deviations—rather than their realized actions or their realized beliefs—is the formal expression of the Lucas critique in this environment: if player $i$ changes their control law, the state dynamics change, which changes every opponent’s signal process, filtering problem, and actions, which feeds back into the state.
This paper aims to establish that the class of noise-state linear strategies is closed under best responses and to derive a deterministic impulse-response system whose fixed points correspond to Nash equilibria in that class.
Remark 2.2 (Subjective equilibrium and policy evaluation).
The best-response closure (Corollary 3.5) does not require player $i$ to know opponents’ maps correctly: a player holding incorrect beliefs about opponents’ strategies still solves a well-posed best-response problem against the believed profile. A policy-maker can therefore let each agent optimize under its own beliefs, then evaluate welfare through the realized state dynamics (2.1), which depend on the actual strategy profile rather than on whether agents’ beliefs about each other are correct. Common knowledge of rationality is sufficient for Nash equilibrium but is not required to apply the characterization as a best-response operator.
Noise-state.
For each player $i$ and each $t\in[0,T]$, define the noise-state as the conditional mean path of the aggregated primitive noise:
$\widehat{W}_{t}^{i}(u):=\mathbb{E}[W_{u}\mid\mathcal{F}_{t}^{i}],\qquad 0\leq u\leq t.$ (2.5)
In linear–Gaussian settings, the conditional law of primitive shocks is Gaussian with deterministic conditional covariance, so $\widehat{W}_{t}^{i}(\cdot)$ is a complete (path-valued) summary of player $i$’s private information.
Definition 2.2 ((Primitive-noise) Linear Process).
An $L^{2}$ stochastic process $L=(L_{t})_{t\in[0,T]}$ is a (primitive-noise) Linear process if there exist deterministic functions
$\bar{L}:[0,T]\to\mathbb{R}^{m},\qquad L:[0,T]^{2}\to\mathbb{R}^{m\times(n+1)d},$
such that, for every $t\in[0,T]$,
$L_{t}\;=\;\bar{L}(t)+\int_{0}^{t}L(t,s)\,dW_{s},\qquad\int_{0}^{t}\|L(t,s)\|^{2}\,ds<\infty.$
When convenient, we extend $L(t,s)$ by 0 for $s>t$.
Remark 2.3 (Impulse-response notation).
We write $X_{t}(s)$ for the deterministic impulse response, not the random variable $X_{t}$ evaluated at an argument; the shock index $s$ distinguishes the two. The same subscript-$t$, parenthetical-shock convention applies to all two-time response maps.
Remark 2.4 (Properties of primitive-noise linear processes).
By construction, primitive-noise linear processes are $\mathbb{F}$-adapted Gaussian processes. Moreover, by Itô isometry the square-integrability condition above is equivalent to $L_{t}\in L^{2}$ for each $t$. Finally, the martingale representation theorem implies that any $L^{2}$ $\mathcal{F}_{t}$-measurable random variable admits a representation of the form $\int_{0}^{t}Z_{s}\,dW_{s}$ with a stochastic integrand $Z$; the defining restriction of the linear class is that the response map $L(t,\cdot)$ is deterministic (for each fixed $t$).
Definition 2.3 (Noise-state Linear control / strategy).
Fix a player $i$ and recall their noise-state path $\widehat{W}_{t}^{i}(\cdot)$ (Section 3). An admissible control $D^{i}=(D_{t}^{i})_{t\in[0,T]}$ is a noise-state linear control if it is affine in the noise state. That is, if there exist deterministic functions
$\bar{D}^{i}:[0,T]\to\mathbb{R}^{d},\qquad D^{i}:[0,T]^{2}\to\mathbb{R}^{d\times(n+1)d},$
such that, for every $t\in[0,T]$,
$D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u),\qquad\int_{0}^{t}\|D^{i}_{t}(u)\|^{2}\,du<\infty,$
and we extend $D^{i}_{t}(u)$ by 0 for $u>t$. Equivalently, writing $\widehat{W}^{\,i}[Y^{i}]$ to emphasize dependence on the observation path, this defines the strategy map
$Y^{i}_{\cdot\wedge t}\longmapsto D_{t}^{i}[Y^{i}]:=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{\,i}[Y^{i}](u).$
3 Main results
We work in the linear–Gaussian model of Section 2. Two results drive the analysis. First, when opponents use noise-state linear strategies, the filtering equations inherit deterministic impulse responses (Theorem 3.1). Second, the best response over the full admissible class is itself noise-state linear (Theorem 3.2), so Nash equilibrium reduces to a deterministic fixed point in impulse-response maps. The filtering closure (Theorem 3.1) shows that noise-state dynamics have deterministic kernels; the best-response closure (Theorem 3.2) derives the adjoint system and the information wedge; Proposition 3.6 then decomposes equilibrium cost into a certainty-equivalent term and an information cost linked to endogenous precision.
Filtering closure.
Recall the noise-state $\widehat{W}_{t}^{i}(\cdot)$ (2.5). We record the deterministic-kernel representation of the estimated-noise increment and its innovations gain in noise-state coordinates. Two-time kernels throughout live on the causal triangle $\Delta_{T}:=\{(t,s)\in[0,T]^{2}:0\leq s\leq t\leq T\}$, extended by 0 off $\Delta_{T}$. We write $d_{u}$ for increments in the path index $u$ at fixed estimation time $t$, and $d_{t}$ for updates in $t$ at fixed $u$. The state is assumed to admit the linear form $X_{t}=\bar{X}_{t}+\int_{0}^{t}X_{t}(s)\,dW_{s}$, where $X_{t}(s)\in\mathbb{R}^{d\times(n+1)d}$ is a deterministic impulse-response kernel mapping primitive shocks into the physical state.
Theorem 3.1 (Linear Filtering Closure).
Fix a player $i$. Assume the state admits the primitive-noise linear form
$X_{t}=\bar{X}_{t}+\int_{0}^{t}X_{t}(s)\,dW_{s},$
and player $i$ observes
$dY_{t}^{i}=\sqrt{P^{i}(t)}\,X_{t}\,dt+E^{i}\,dW_{t},$
with deterministic $P^{i}(t)\succeq 0$ and block selector $E^{i}$.
Let $\Pi^{i}:=E^{i,\top}E^{i}$ and $\widehat{X}_{t}^{i}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{i}]$, and define the innovation $dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$.
(i) The estimated-noise path increments admit the decomposition
$d_{u}\widehat{W}_{t}^{i}(u)=\Pi^{i}\,dW_{u}+\Big(\int_{0}^{t}F_{t}^{i}(u,s)\,dW_{s}\Big)\,du,\qquad u<t,$ (3.1)
where the deterministic kernel $F^{i}$ is given explicitly by
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (3.2)
(ii) The mixed $(t,u)$-update satisfies
$d_{t}\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du.$
The unresolved kernel $\widetilde{X}^{i}$ is determined algebraically from $F^{i}$ by (A.12); the filtering kernel $F^{i}$ is the unique solution of the forward evolution system (A.14)–(A.15) (Definition A.1).
Remark 3.1 (Regression interpretation and scale separation).
Discretizing time, each observation increment is affine in the noise increments $\Delta W_{1:k}$, so by joint Gaussianity the conditional mean $\mathbb{E}[\Delta W_{1:k}\mid\Delta Y^{i}_{1:k}]$ is a deterministic linear projection; each layer of “beliefs about beliefs” is a further deterministic composition of projections, which is why the regress closes. The decomposition in part (i) reflects a scale separation: the direct component $\Pi^{i}\,dW_{u}$ is $O(1)$ (player $i$ observes their own noise contemporaneously), while the indirect component—absolutely continuous in $u$ with density driven by $F^{i}$—captures drift-based inference about unobserved noise coordinates at scale $O(dt)$.
Best-response closure and information wedges.
When opponents use noise-state linear strategies, a player’s best response over the full admissible $L^{2}$ class remains noise-state linear; the maximum principle yields a closed backward system of deterministic adjoint kernels including the information wedge $\mathcal{V}^{i}(t)$ that prices bilateral belief manipulation.
Theorem 3.2 (Multiplayer belief-adjoint kernels and information wedges).
Fix a player $i$ and suppose each opponent $k\neq i$ uses a noise-state linear strategy with deterministic impulse-response maps, which pins down a deterministic forward environment (state impulse response $X(\cdot,\cdot)$, filtering objects $(P^{k},\widetilde{X}^{k},F^{k})$, and policy kernels $D^{k}$).
Structure. Because the equilibrium control is a noise-state linear process $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,dW_{u}$ with deterministic mean and kernel, the backward system splits into two structurally identical subsystems—one for the mean coefficients $(\bar{H}^{X,i},\bar{H}^{k,i})$, one for the response-map coefficients $(H^{X,i},H^{k,i})$—that should be read in parallel.
In each subsystem, the single-agent costate acquires an additional coupling term: player $i$’s action shifts the state, which shifts opponent $k$’s signal, posterior, and action, feeding back into the state. This channel introduces belief-adjoint variables $(\bar{H}^{k,i},H^{k,i})$ that track the shadow value of perturbing opponent $k$’s noise-state estimate, coupled into the physical costate through the opponents’ filtering gains.
Mean backward system. The mean physical costate $\bar{H}^{X,i}_{t}$ and mean belief-adjoints $\bar{H}^{k,i}_{t}(u)$ satisfy
$\frac{d}{dt}\bar{H}^{X,i}_{t}$ $=-\big(G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\big)-A(t)^{\top}\bar{H}^{X,i}_{t}-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz}_{=:\;\bar{\mathcal{V}}^{i}_{t},\;\text{mean information wedge}},$ (3.3)
$\frac{d}{dt}\bar{H}^{k,i}_{t}(u)$ $=-\big(D^{k}_{t}(u)\big)^{\top}\bar{H}^{X,i}_{t}+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz,\qquad k\neq i,$ (3.4)
with terminal conditions $\bar{H}^{X,i}_{T}=G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}$ and $\bar{H}^{k,i}_{T}(\cdot)=0$.
The first two terms in (3.3) are the standard single-agent costate equation; the third term $\bar{\mathcal{V}}^{i}_{t}$ is the mean information wedge, absent in the single-agent problem. In the belief-adjoint equation (3.4), the forcing $D^{k}_{t}(u)^{\top}\bar{H}^{X,i}_{t}$ transmits the physical costate into the belief channel through opponent $k$’s policy kernel: the larger $D^{k}_{t}(u)$, the more opponent $k$’s action at time $t$ responds to the $u$-th shock, and the more valuable it is for player $i$ to have shifted $k$’s estimate of that shock. The integral term is a self-coupling through which belief perturbations propagate forward via the filtering gain $\widetilde{X}^{k}$.
Kernel backward system. The kernel costate $H^{X,i}_{t}(r)$ and kernel belief-adjoints $H^{k,i}_{t}(u,r)$ satisfy the identical system with $\bar{X}_{t}\to X(t,r)$ and $\bar{G}^{X,i}_{t}\to G^{X,i}_{t}(r)$; the kernel information wedge $\mathcal{V}^{i}_{t}(r)$ is the corresponding substitution of $\bar{\mathcal{V}}^{i}_{t}$. We define the (full) information wedge as the linear process
$\mathcal{V}^{i}_{t}:=\bar{\mathcal{V}}^{i}_{t}+\int_{0}^{t}\mathcal{V}^{i}_{t}(r)\,dW_{r},$ (3.5)
which is the total distortion of the physical costate relative to the single-agent benchmark. The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ captures how belief manipulation distorts the equilibrium level (mean actions); the response-map wedge $\mathcal{V}^{i}_{t}(r)$ captures how it distorts the equilibrium response to shocks. Both components—and hence $\mathcal{V}^{i}_{t}$ itself—vanish when signals are exogenous.
Policy interpretation of the wedge.
The information wedge $\mathcal{V}^{i}_{t}$ vanishes identically when signals are exogenous to controls, recovering the standard LQG costate equation. When controls enter the state dynamics, a change in signal precision $P^{k}(t)$ also operates through the belief-manipulation channel: it alters how sharply opponents’ posteriors respond to drift perturbations (through $\widetilde{X}^{k}$), which alters the shadow value of shifting those posteriors (through $\bar{H}^{k,i}$), which feeds back into the equilibrium control through (3.3). Any model that omits this channel misattributes the strategic distortion to the physical response: the mean error is exactly $\bar{\mathcal{V}}^{i}_{t}$, and the shock-by-shock error is $\mathcal{V}^{i}_{t}(r)$. The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ governs the distortion in equilibrium mean actions (Section 4.3); the kernel wedge $\mathcal{V}^{i}_{t}(r)$ governs the distortion in equilibrium impulse responses. Section 4 quantifies both channels numerically.
Corollary 3.3 (Exogenous-signal reduction).
If controls do not enter the state dynamics, then $\mathcal{V}^{i}_{t}\equiv 0$ for all $i$ and $t$, the belief-adjoint equations (3.4) decouple, and the equilibrium reduces to $n$ independent single-agent LQG problems against a common exogenous state.
Lemma 3.4 (Strict convexity of the best-response problem).
Fix opponents’ strategy maps $D^{-i}$ in the deterministic impulse-response class (hence fixed linear operators from signal histories to actions). Then player $i$’s induced objective $D^{i}\mapsto J^{i}(D^{i};D^{-i})$ is a strictly convex functional on the admissible $L^{2}$ control space. Consequently, the spike-variation stationarity condition (B.8) is not only necessary but sufficient: any admissible control satisfying it is the unique global best response.
Proof sketch.
With opponents’ maps fixed and linear, the closed-loop drift of $X$ is affine in $D^{i}$, so $(X,D^{-i})$ depend affinely on $D^{i}$. The cost (2.4) is quadratic with $G^{DD,i}(t)\succ 0$, hence strictly convex in $D^{i}$. ∎
Corollary 3.5 (Noise-state linear best response (global closure)).
Under the conditions of Theorem 3.2, fix an opponents’ profile $D^{-i}$ of noise-state linear strategies (Definition 2.3). If $D^{i}$ is a best response to $D^{-i}$ over the full admissible $L^{2}$ control class, then for a.e. $t\in[0,T]$,
$G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u)=0,$
and hence
$D_{t}^{i}=-(G^{DD,i}(t))^{-1}\bar{H}^{X,i}_{t}-(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u).$
In particular, every best response over the full admissible class is itself a noise-state linear control.
Proof.
With opponents’ maps fixed and linear, the innovation $dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion whose law does not depend on $D^{i}$ (Liptser–Shiryaev [16], Theorem 7.12). Since $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$ (Theorem 3.1 in innovations coordinates), player $i$’s control is a deterministic affine functional of the exogenous process $I^{i}$. The induced objective is strictly convex (Lemma 3.4), so the stationarity condition (B.8) is necessary and sufficient, yielding the stated formula. Note that the exogeneity of $I^{i}$ is a property of the best-response problem: the response-maps mediating between $I^{i}$ and the noise-state depend on the full strategy profile through the forward closure, so the equilibrium remains a genuine fixed point (Theorem C.5). ∎
Remark 3.2.
Since every best response over the full admissible $L^{2}$ class is noise-state linear (Corollary 3.5), any Nash equilibrium of the restricted game in which players are confined to noise-state linear controls is automatically a Nash equilibrium of the unrestricted game. The open question is whether equilibria outside the deterministic impulse-response class exist; Section 6 discusses a program for ruling them out.
When $n=1$, $\mathcal{V}^{i}_{t}\equiv 0$ and the ansatz $H^{X,i}_{t}(r)=S(t)X(t,r)$ reduces the backward system to the standard matrix Riccati equation for $S$, recovering classical separated LQG control.
Risk-sensitive extension.
Replace (2.4) with the entropic risk measure $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$, which penalizes cost variance in addition to cost mean ($\theta_{i}>0$ governs risk aversion; $\theta_{i}\to 0$ recovers (2.4)). The entire framework carries over: the best-response FOC is identical to Corollary 3.5 with the noise-state $\widehat{W}_{t}^{i}$ replaced by a risk-adjusted noise-state $\widehat{W}^{\,i,\theta}_{t}$ that tilts the conditional mean of $W$ toward directions that reduce expected cost and rescales conditional precision along directions where cost variance is large. The filtering closure is unchanged; certainty equivalence fails because the quadratic structure of the cost reshapes the posterior without altering the observation channel. The full statement and derivation are in Appendix B.6 (Theorem B.1).
Information cost and endogenous precision.
At the equilibrium profile, the control gap between full and partial information is
$D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$ (3.6)
where $D^{i,\mathrm{full}}_{t}:=-(G^{DD,i}(t))^{-1}M^{i}_{t}$ is the best response under full observation of $W$, $M^{i}_{t}:=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}W_{u}$ is the stochastic costate, and $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$ is the noise-state estimation error.
Proposition 3.6 (Information cost).
At the equilibrium profile,
$\Delta J^{i}:=J^{i}(D^{i};D^{*,-i})-J^{i}(D^{i,\mathrm{full}};D^{*,-i})=\int_{0}^{T}\mathrm{tr}\!\left[G^{DD,i}(t)^{-1}\,\Sigma^{M,i}_{t}\right]\mathrm{d}t,$ (3.7)
where $\Sigma^{M,i}_{t}:=\operatorname{Cov}(M^{i}_{t}\mid\mathcal{F}^{i}_{t})$ is the costate estimation error covariance (deterministic by Gaussianity). Moreover, $\Delta J^{i}$ depends on $P^{i}$ only through the filtering kernels $(\widetilde{X}^{i},F^{i})$ and is independent of player $i$’s control $D^{i}$.
The total cost decomposes as
$J^{i}(D^{i};D^{*,-i})=\underbrace{J^{i}(D^{i,\mathrm{full}};D^{*,-i})}_{\text{certainty-equivalent cost}}+\;\underbrace{\Delta J^{i}(P^{i})}_{\text{information cost}},$ (3.8)
This decomposition underlies the welfare comparisons in Section 4: the information wedge enters through $H^{X,i}$ inside $\Sigma^{M,i}_{t}$. The joint optimization over $(D^{i},P^{i})$ separates: $D^{i}$ is determined by Corollary 3.5 for any given $P^{i}$, and $P^{i}$ solves
$\min_{P^{i}(\cdot)\succeq 0}\quad\Delta J^{i}(P^{i})+\int_{0}^{T}c^{i}(t,P^{i}(t))\,\mathrm{d}t,$ (3.9)
where $c^{i}$ is a convex attention cost.
Corollary 3.7 (Endogenous precision closure).
If each opponent $k\neq i$ uses a deterministic precision path $P^{k}(\cdot)$, then $\Delta J^{i}$ is a deterministic functional of $P^{i}(\cdot)$, and the best-response precision is deterministic.
Proof.
Immediate from Proposition 3.6: the forward environment and backward adjoint are deterministic, so (3.9) is a deterministic variational problem. ∎
Remark 3.3 (Connection to rational inattention).
When $n=1$ the FOC for (3.9) reduces to the continuous-time analogue of Sims [27]; with $n>1$ the information wedge enters through $H^{X,i}$, so the marginal value of precision depends on the strategic environment—a more responsive opponent raises the return to filtering because better inference enables more effective belief manipulation.
Remark 3.4 (Infinite horizon and the frequency domain).
Under time-homogeneous primitives, response maps become time-invariant K(t,u)=k(tu)K(t,u)=k(t-u) and the fixed point becomes one in transfer functions, extending [12, 14, 24] to endogenous signals. Rational transfer functions correspond to finite-dimensional Markov sufficient statistics, so the pole structure of equilibrium kernels diagnoses whether an exact state-space reduction exists.
Reference: impulse-response objects.
For convenience, we collect the kernel objects introduced above.
$W=(W^{0},\dots,W^{n})$ primitive noise $E^{i},\;\Pi^{i}:=E^{i,\top}E^{i}$ selector / projector
$\widehat{W}_{t}^{i}(u)$ noise-state $X_{t}(s)$ state impulse response
$F_{t}^{i}(u,s)$ filtering impulse response $\widetilde{X}^{i}_{t}(u)$ unresolved state response
$D^{i}_{t}(u)$ policy impulse response $\mathcal{D}^{i}_{t}(s)$ primitive-noise control
$(\bar{H}^{X,i},H^{X,i})$ physical adjoints $(\bar{H}^{k,i},H^{k,i})$ belief adjoints
$\mathcal{V}^{i}_{t}=\bar{\mathcal{V}}^{i}_{t}+\int_{0}^{t}\mathcal{V}^{i}_{t}(r)\,dW_{r}$ information wedge
4 Bilateral belief manipulation in a two-player game
To make the general theory concrete, we specialize to the simplest game that exhibits bilateral belief manipulation: two symmetric players with scalar state, opposing targets, and equal signal precision. This strips away all structure except the core strategic friction. Neither player observes the state directly; each receives a private diffusion signal of precision $p>0$. Because each player’s control enters the state dynamics, it reshapes the opponent’s signal and thereby the opponent’s posterior. The information wedge $\mathcal{V}^{i}_{t}$ of Theorem 3.2 prices this externality exactly; the mean wedge $\bar{\mathcal{V}}^{i}_{t}$ is the component we visualize below.
4.1 Primitives, objectives, and symmetry
Fix a horizon $[0,T]$. The state $X_{t}\in\mathbb{R}$ satisfies
$dX_{t}=\big(D_{t}^{1}+D_{t}^{2}\big)\,dt+dW_{t}^{0},\qquad X_{0}=x_{0},$ (4.1)
where $W^{0}$ is a standard Brownian motion. Player $i\in\{1,2\}$ does not observe $X$ directly; instead they observe the diffusion channel
$dY_{t}^{1}=\sqrt{p}X_{t}\,dt+dW_{t}^{1},\qquad dY_{t}^{2}=\sqrt{p}X_{t}\,dt+dW_{t}^{2},$ (4.2)
with $(W^{0},W^{1},W^{2})$ mutually independent and $p>0$ representing root precision. Player $i$ chooses $D^{i}$ adapted to $\mathcal{F}_{t}^{i}=\sigma(Y_{s}^{i}:s\leq t)$.
Running costs are symmetric tracking losses with quadratic effort penalty $r>0$:
$J^{1}(D^{1};D^{2})=\mathbb{E}\!\int_{0}^{T}\Big((X_{t}-1)^{2}+r\,(D_{t}^{1})^{2}\Big)\,dt,\quad J^{2}(D^{2};D^{1})=\mathbb{E}\!\int_{0}^{T}\Big((X_{t}+1)^{2}+r\,(D_{t}^{2})^{2}\Big)\,dt.$ (4.3)
There is no terminal cost in this worked example.
The model is invariant under the sign/label symmetry
$(X,\;Y^{1},\;Y^{2},\;D^{1},\;D^{2})\ \longmapsto\ (-X,\;Y^{2},\;Y^{1},\;-D^{2},\;-D^{1}).$
[…]
Figure 2: Information-to-action feedback in the worked example. Actions move $X_{t}$ and shift opponents’ beliefs, so equilibrium mean actions depend on precision $p$.
4.2 Numerical equilibrium and kernel visualization
We now present the numerical equilibrium obtained from the Picard iteration described above. All figures below are generated from the converged kernel profile $D^{1}$, its primitive-noise representation $\mathcal{D}^{1}$, the state kernel $X$, and the filtering kernel $F^{1}$.
Computation.
Equilibrium is computed by iterating the best-response operator on the deterministic impulse-response system, discretized on a uniform grid of $N=40$ points; no Monte Carlo simulation is needed. At each iterate the forward environment $(X,\widetilde{X}^{i},F^{i})$ is recomputed from the closure equations, the backward adjoint $H^{X,i}$ is solved by explicit backward Euler, and the updated policy is blended into the current profile. Parameters: $T=1$, $x_{0}=0$, $r=0.1$, $G^{XX,i}(T)=0$. The outer residual reaches $10^{-5}$ in $19$ iterations( ${\approx}\,10$ ms in C++). At longer horizons the binding constraint is representational (the filtering kernel becomes more singular) rather than economic.
Convergence of the fixed point.
The outer Picard residual decays geometrically, confirming convergence of the fixed-point scheme in this symmetric benchmark; we use the iteration as a constructive computation rather than an existence proof.
Impulse-response viewpoint.
The key distinction is between $D^{1}_{t}(s)$ (response to a shock as inferred through the filter).
(a) Primitive-noise state kernel $X_{t}(s)$.
(b) Equilibrium noise-state feedback kernel $D^{1}_{t}(s)$.
Figure 3: Equilibrium kernel time-slices, shown channel-by-channel for $(W^{0},W^{1},W^{2})$ with curves colored by evaluation time $t$. (a) Both players jointly stabilize older fundamental shocks (decay in the $W^{0}$ panel); observation-noise channels show smaller, opposed effects mediated by the belief loop. (b) The system-noise channel dominates; observation-noise channels exhibit curvature reflecting drift-based inference.
4.3 Mean actions, separation failure, and bilateral belief manipulation
Since the stochastic integral in $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$ has zero unconditional mean, $\mathbb{E}[D_{t}^{i}]=\bar{D}^{i}_{t}$ isolates the deterministic tug-of-war over the target $\pm 1$, while the kernel term captures impulse responses to posterior shocks. Under perfect information, the mean policy does not depend on signal precision—a certainty-equivalence property. Decentralized information breaks this: changing precision changes the speed at which opponents update, giving each player a reason to act on posteriors rather than just on the physical state.
Figure 4: Equilibrium mean control $\bar{D}^{1}_{t}$ as a function of signal precision $p$, with the perfect-information benchmark (dashed). Low precision makes opponents’ posteriors sluggish, amplifying the incentive to manipulate beliefs; as $p\to\infty$ the mean policy converges to the perfect-information limit.
The information wedge as a dynamic rent density.
In the general theory (Theorem 3.2), the backward equation for $\bar{H}^{X}$ contains opponent-filter coupling terms that vanish under perfect information or in the single-player case. Here these terms aggregate into
$\bar{\mathcal{V}}^{1}_{t}=P^{2}(t)\int_{0}^{t}\widetilde{X}^{2}_{t}(z)\cdot\bar{H}^{2}_{t}(z)\,dz,$
which enters the mean adjoint equation as $\tfrac{d}{dt}\bar{H}^{X}_{t}=-(\bar{X}_{t}-1)-\bar{\mathcal{V}}^{1}_{t}$. The unresolved kernel $\widetilde{X}^{2}_{t}(z)$ measures how much of the $z$-th shock remains unresolved by player 2 at time $t$; the belief-adjoint $\bar{H}^{2}_{t}(z)$ measures the shadow value to player 1 of that shift. Their inner product $\bar{\mathcal{V}}^{1}_{t}$ prices the mean channel of bilateral belief manipulation at time $t$; Figures 56 visualize this object across precision asymmetries.
Figure 5: Asymmetric equilibrium with $p_{1}=3$ fixed and $p_{2}$ varying. Left: mean controls $\bar{D}^{1}_{t}$ (solid) and $\bar{D}^{2}_{t}$ (dashed) diverge as the precision gap widens. Center: the mean state path $\bar{X}_{t}$ tilts toward the better-informed player’s target. Right: aggregate mean effort $|\bar{D}^{1}|+|\bar{D}^{2}|$ exceeds the perfect-information benchmark (gray dashed), with the excess growing in the precision asymmetry.
Figure 6: Mean information wedges $\bar{\mathcal{V}}^{1}_{t}$ and $\bar{\mathcal{V}}^{2}_{t}$ as $p_{2}$ varies ($p_{1}=3$ fixed). Both wedges are hump-shaped in $t$, confirming that belief manipulation is most valuable at intermediate dates. Increasing the opponent’s precision amplifies the wedge: a better-informed opponent reacts more sharply to drift, raising the marginal value of manipulating their posterior. The wedges vanish at the terminal date, consistent with the zero terminal condition on the belief adjoint.
Figure 7: Equilibrium costs under private signals (dashed) and pooled signals (solid) with $p_{1}=3$ fixed and $p_{2}$ varying. Left: competitive targets ($\theta_{1}=1$, $\theta_{2}=-1$). Right: cooperative targets ($\theta_{1}=\theta_{2}=0$). Pooling is Pareto-improving in both cases, but the gain is an order of magnitude larger under competition, where it eliminates the bilateral belief-manipulation externality priced by $\mathcal{V}^{i}_{t}$. Blue: player 1; red: player 2.
Information pooling as a Pigouvian intervention.
Figure 7 compares equilibrium costs under private signals with those under a pooled signal of precision $P^{1}+P^{2}$, for both competitive ($\theta_{1}=1$, $\theta_{2}=-1$) and cooperative ($\theta_{1}=\theta_{2}=0$) targets. Pooling is Pareto-improving in both cases, but the gains differ by an order of magnitude: roughly 0.020.020.050.05 under cooperation versus 0.30.30.50.5 under competition. Since the estimation benefit of pooling enters identically regardless of the target structure, the difference isolates the strategic externality channel priced by $\mathcal{V}^{i}_{t}$: nearly all of the welfare gain from mandatory disclosure comes from eliminating bilateral belief manipulation, not from improving state estimation.
5 Market microstructure as a special case
The two-player example operates within the quadratic-cost framework of Section 2. We now show that the noise-state technique embeds the Kyle–Back model with no additional machinery, despite traders maximizing expected profits against the market maker’s price rather than minimizing a quadratic objective. The embedding works because $\lambda(t)>0$ provides the strict convexity needed for the best-response closure, the filtering closure applies verbatim to the stacked observation channel, and the backward system of Appendix D yields the optimality condition. We derive the equilibrium characterization for a single informed trader and identify the two channels through which belief manipulation operates; the extension to multiple asymmetrically informed traders follows the same backward system and is a natural application of the general theory.
5.1 Model and embedding
Adjoin a fundamental-value Brownian motion $W^{V}$ independent of $(W^{0},\dots,W^{n})$, with selector $E^{V}$ and projector $\Pi^{V}$. The fundamental is $V_{t}:=\sigma_{V}W_{t}^{V}$. All players observe aggregate order flow $dZ_{t}=\sum_{j}D_{t}^{j}\,dt+\sigma_{0}\,dW_{t}^{0}$; trader $i$ additionally observes a private signal $dS_{t}^{i}=\sqrt{p_{i}}\,V_{t}\,dt+dW_{t}^{i}$. The market maker observes only $Z$ and posts the competitive price $P_{t}:=\mathbb{E}[V_{T}\mid\mathcal{F}_{t}^{0}]=\sigma_{V}\int_{0}^{t}E^{V}d_{u}\widehat{W}_{t}^{0}(u)$. Trader $i$ maximizes $\pi^{i}:=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$. By symmetry $W\mapsto-W$, all mean terms vanish.
Embedding.
Rescaling the order-flow channel to unit noise, trader $i$’s stacked observation is
$d\!\begin{pmatrix}Z_{t}/\sigma_{0}\\ S_{t}^{i}\end{pmatrix}=\underbrace{\begin{pmatrix}\sigma_{0}^{-1}\\[2.0pt] \sqrt{p_{i}}\end{pmatrix}}_{\sqrt{P^{i}}}\!\!\begin{pmatrix}\text{demand drift}\\ \sigma_{V}E^{V}W_{t}\end{pmatrix}\!dt\;+\;\underbrace{\begin{pmatrix}E^{0}\\ E^{i}\end{pmatrix}}_{\mathcal{E}^{i}}\!dW_{t},$ (5.1)
which is the observation equation of Section 2 with $\sqrt{P^{i}}=(\sigma_{0}^{-1},\sqrt{p_{i}})^{\top}$ and $E^{i}\to\mathcal{E}^{i}$; the market maker has $\sqrt{P^{0}}=\sigma_{0}^{-1}$, $\mathcal{E}^{0}=E^{0}$.
Kyle’s lambda.
Define the aggregate primitive-noise demand kernel $\mathcal{D}_{\Sigma,t}(s):=\sum_{j}\mathcal{D}^{j}(t,s)$. Since $E^{V}E^{0,\top}=0$, all price information comes through drift-based inference:
$dP_{t}=\lambda(t)\,dI_{t}^{0},\qquad\lambda(t):=\sigma_{V}\!\int_{0}^{t}E^{V}\widetilde{C}^{0}(t,u)^{\top}du,$ (5.2)
where $\widetilde{C}^{0}$ is the market maker’s unresolved drift kernel under the projection decomposition (A.12).
Remark 5.1 (Quadratic structure of the profit).
Fix opponents’ strategy maps. Since the market maker’s filtering gains are deterministic, the price decomposes as $P_{t}=P_{t}^{-i}+\int_{0}^{t}\ell(t,s)\,D_{s}^{i}\,ds$, where $P_{t}^{-i}$ is a linear process independent of $D^{i}$ and $\ell(t,s)$ is a deterministic kernel. Symmetrically, the future execution cost of a unit impulse at time $t$ is $\int_{t}^{T}D_{\tau}^{i}\,\delta P_{\tau}\,d\tau$, where $\delta P_{\tau}$ propagates forward through the deviation system with deterministic coefficients. The profit $\pi^{i}=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$ is therefore concave quadratic in $D^{i}$, when $\lambda(t)>0$ provides strict concavity.
5.2 Equilibrium characterization
The optimality condition from a unit demand impulse at $t$ equates the conditional mispricing to the integrated impact cost (Appendix D):
$\underbrace{\sigma_{V}E^{V}\,\partial_{u}\bigl(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0}\bigr)(u)}_{H^{i,\mathrm{val}}_{t}(u):\;\text{information rent density}}\;=\;\underbrace{\int_{t}^{T}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau}_{H^{i,\mathrm{imp}}_{t}(u):\;\text{impact cost density}},$ (5.3)
where $\delta P_{\tau}$ is the price perturbation for an impulse at $t$. Differentiating both sides in $t$ and solving for $D^{i}_{t}(u)$ yields a closed backward ODE:
$\lambda(t)\,D^{i}_{t}(u)=-\frac{d}{dt}H^{i,\mathrm{val}}_{t}(u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau,$ (5.4)
with terminal condition $D^{i}_{T}(u)=\lambda(T)^{-1}H^{i,\mathrm{val}}_{T}(u)$.
Assumption 5.1 (Persistent price informativeness).
$\lambda(t)>0$ for all $t\in[0,T]$.
This requires that aggregate order flow remains informative about the fundamental at every date, the economically relevant case since temporary uninformativeness would present an arbitrage opportunity. The condition is self-reinforcing: positive $\lambda$ induces finite trading rates, which maintain the drift informativeness that sustains $\lambda>0$. It plays the role of $G^{DD,i}\succ 0$ (Lemma 3.4), ensuring strict convexity of the best-response problem on the domain of the fixed-point operator.
The derivative $\frac{d}{dt}H^{i,\mathrm{val}}$ and the deviation system determining $\delta P_{\tau}$ are given in Appendix D.
5.3 Two channels of belief manipulation
Decompose $\delta P_{\tau}$ into $\delta P_{\tau}^{\mathrm{dir}}$ (computed with $D^{k}\equiv 0$, i.e. the market maker’s response alone) and $\delta P_{\tau}^{\mathrm{ind}}:=\delta P_{\tau}-\delta P_{\tau}^{\mathrm{dir}}$ (opponent-mediated). Since $\delta P_{\tau}^{\mathrm{dir}}=\lambda(t)$ for all $\tau\geq t$, the impact adjoint splits into:
  1. (i)
    $\mathcal{V}^{i,\mathrm{price}}(t,u)=\lambda(t)\int_{t}^{T}D^{i}_{\tau}(u)\,d\tau$: instantaneous impact times the integrated future policy kernel. This is the standard Kyle channel—the trader internalizes how each demand impulse moves the price.
  2. (ii)
    $\mathcal{V}^{i,\mathrm{opp}}(t,u)=\int_{t}^{T}\delta P_{\tau}^{\mathrm{ind}}\,D^{i}_{\tau}(u)\,d\tau$: opponent-mediated impact, arising because opponents also observe order flow and adjust their trading, which feeds back into prices.
Back’s model.
When $n=1$ and $p_{1}=0$ (Back’s model), the opponent channel vanishes ($\delta P_{\tau}^{\mathrm{ind}}\equiv 0$) and (5.3) reduces to Back’s integral equation for the optimal trading rate. With $n\geq 2$ and $p_{i}>0$ both channels are active: each trader tempers demand not only to limit direct price impact but to prevent opponents from inferring and front-running through shared order flow. Guess-and-verify cannot close this system; the noise-state technique handles it with no additional machinery.
6 Conclusion
Conditioning on primitive Brownian shocks collapses the Townsend belief hierarchy onto deterministic impulse-response maps, yielding the first exact equilibrium characterization of finite-player continuous-time LQG games with endogenous signals. The information wedge $\mathcal{V}^{i}_{t}$ prices bilateral belief manipulation and vanishes when signals are exogenous. The two-player example shows that nearly all welfare gains from disclosure come from eliminating this channel; the Kyle–Back embedding shows the technique extends beyond quadratic costs, identifying a second channel through opponent inference via shared order flow that applies to asymmetric precisions and continuous private learning.
Further directions.
The main open question is whether the deterministic impulse-response class is without loss of generality. A natural program proceeds in three steps: (i) a Blackwell reduction showing that under quadratic costs the best response depends on the posterior $\pi_{t}^{i}:=\mathcal{L}(W_{[0,t]}\mid\mathcal{F}_{t}^{i})$ only through its mean and covariance; (ii) a belief-of-belief collapse in which player $i$’s estimate of $\pi_{t}^{k}$ factors through a deterministic operator—the posterior analogue of the kernel maps $(F^{i},\widetilde{X}^{i})$; (iii) a contraction argument on the variance functional $V(\mathbf{D}):=\sum_{i}\mathbb{E}\int\!\!\int\|D^{i}(t,s,\pi_{t}^{i})-\mathbb{E}[D^{i}(t,s,\pi_{t}^{i})]\|^{2}\,ds\,dt$, which vanishes exactly in the deterministic impulse-response class and which iterated best responses are expected to contract. Beyond this, stochastic coefficients (Remark C.2), graph-structured interactions, and global uniqueness via the dissipative structure of $F^{i}$ are natural extensions.
Appendix A Filtering
Roadmap.
Throughout, formal calculations identify the kernel equations for $F^{i}$; we then define $F^{i}$ as the unique solution of those equations, with uniqueness following from uniqueness of $L^{2}$-projection. The pipeline proceeds in three steps. A.1 derives the dynamics of $t\mapsto\widehat{W}_{t}^{i}(u)$ from a characterizing identity for conditional expectations, using an ansatz with undetermined coefficients matched against exponential test martingales. A.2 specializes the abstract dynamics to the physical observation model, identifying $\widetilde{X}^{i}_{t}(u)$ as the filtering gain via Itô isometry, establishing the algebraic identity $\widetilde{X}^{i}=X(I-\Pi^{i})-X*F^{i}$, and recording the contraction identity $\int_{0}^{t}X_{t}(s)\,\widetilde{X}^{i}_{t}(s)^{\top}\,ds=\Sigma^{XX,i}(t)$. A.3 defines $F^{i}$ as the primitive filtering object through its forward evolution equation, derives the explicit three-term formula, and verifies uniqueness.
A.1 Dynamics of the noise-state
Fix a player $i$. Recall the noise-state $\widehat{W}_{t}^{i}(u):=\mathbb{E}[W_{u}\mid\mathcal{F}_{t}^{i}]$, $0\leq u\leq t$, and the observation channel
$dY_{t}^{i}=\Big(\bar{C}(t)+\int_{0}^{t}C(t,s)\,dW_{s}\Big)\,dt+E^{i}\,dW_{t},\qquad Y_{0}^{i}=0,$
where $C(t,s)$ is deterministic. Since $\bar{C}$ is deterministic, we may assume without loss that $\bar{C}\equiv 0$ by replacing $Y_{t}^{i}$ with $Y_{t}^{i}-\int_{0}^{t}\bar{C}(r)\,dr$.
Proposition A.1 (Linear closure for $\widehat{W}$).
The map $u\mapsto\widehat{W}_{t}^{i}(u)$ is a linear process: there exists a deterministic indirect-response map $F_{t}^{i}(\cdot,\cdot)$ such that
$\widehat{W}_{t}^{i}(u)=\Pi^{i}W_{u}+\int_{0}^{t}\!\int_{0}^{u}F_{t}^{i}(z,s)\,dz\,dW_{s},\qquad 0\leq u\leq t.$
Proposition A.2 (Dynamics of $\widehat{W}_{t}^{i}(u)$).
There exist deterministic matrix-valued functions $B^{Y}(u,t)$ and $B(u;t,s)$ such that, for each fixed $u$,
$\widehat{W}_{t}^{i}(u)-\widehat{W}_{t_{0}}^{i}(u)=\int_{t_{0}}^{t}\!\!\int_{0}^{r}B(u;r,s)\,d\widehat{W}_{r}^{i}(s)\,dr\;+\;\int_{t_{0}}^{t}B^{Y}(u,r)\,dY_{r}^{i},\qquad t\geq t_{0}\geq u,$ (A.1)
or equivalently in differential form,
$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}B(u;t,s)\,d\widehat{W}_{t}^{i}(s)\Big)\,dt+B^{Y}(u,t)\,dY_{t}^{i},$ (A.2)
with
$B^{Y}(u,t)$ $=\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,C(t,s)^{\top}\,ds\;+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}},$ (A.3)
$B(u;t,s)$ $=-B^{Y}(u,t)\,C(t,s),$ (A.4)
where $\mathsf{C}_{t}^{i}(u,s):=\operatorname{Cov}(W_{u},W_{s}\mid\mathcal{F}_{t}^{i})$ is the deterministic conditional error covariance kernel.
Proof.
We use the characterizing identity for conditional expectations via exponential test martingales. Fix a deterministic vector function $\vec{r}(\cdot)$ and define $\varphi_{t}^{r}$ by $\varphi_{0}^{r}=1$ and $d\varphi_{t}^{r}=i\,\varphi_{t}^{r}\,(dY_{t}^{i})^{\top}\vec{r}(t)$. A necessary and sufficient condition characterizing $\widehat{W}_{t}^{i}(u)$ is
$\mathbb{E}\!\big[\widehat{W}_{t}^{i}(u)\,\varphi_{t}^{r}\big]=\mathbb{E}\!\big[W_{u}\,\varphi_{t}^{r}\big]\qquad\text{for all }\vec{r}(\cdot).$ (A.5)
We compute the time derivative of each side independently, then match.
Left-hand side. Apply Itô’s product rule to $\widehat{W}_{t}^{i}(u)\,\varphi_{t}^{r}$, using the ansatz (A.2) and $d\varphi_{t}^{r}=i\,\varphi_{t}^{r}(dY_{t}^{i})^{\top}\vec{r}(t)$:
$d\big(\widehat{W}_{t}^{i}(u)\,\varphi_{t}^{r}\big)=i\,\widehat{W}_{t}^{i}(u)\,\varphi_{t}^{r}\,(dY_{t}^{i})^{\top}\vec{r}(t)+\varphi_{t}^{r}\,d_{t}\widehat{W}_{t}^{i}(u)+d\langle\widehat{W}^{i}(u),\varphi^{r}\rangle_{t}.$
The quadratic covariation contributes $d\langle\widehat{W}^{i}(u),\varphi^{r}\rangle_{t}=i\,\varphi_{t}^{r}\,B^{Y}(u,t)\,\vec{r}(t)\,dt$. Taking expectations and collecting the $dt$ terms gives
$\frac{d}{dt}\mathbb{E}\!\big[\widehat{W}_{t}^{i}(u)\,\varphi_{t}^{r}\big]$ $=i\,\mathbb{E}\!\bigg[\varphi_{t}^{r}\Big(\widehat{W}_{t}^{i}(u)\!\int_{0}^{t}C(t,s)\,d\widehat{W}_{t}^{i}(s)^{\top}+B^{Y}(u,t)\Big)\bigg]\vec{r}(t)$
$\quad+\mathbb{E}\!\bigg[\varphi_{t}^{r}\int_{0}^{t}\!\big(B(u;t,s)+B^{Y}(u,t)\,C(t,s)\big)\,d\widehat{W}_{t}^{i}(s)\bigg].$ (A.6)
Right-hand side. Expand $\mathbb{E}[W_{u}\,\varphi_{t}^{r}]$ using the SDE for $\varphi^{r}$ and apply the tower property in the form $\mathbb{E}[\varphi_{s}^{r}\,L]=\mathbb{E}[\varphi_{s}^{r}\,\mathbb{E}[L\mid\mathcal{F}_{s}^{i}]]$. Define the conditional second-moment kernel
$\mathcal{M}_{s}^{i}(u,z):=\mathbb{E}[W_{u}W_{z}^{\top}\mid\mathcal{F}_{s}^{i}]=\mathsf{C}_{s}^{i}(u,z)+\widehat{W}_{s}^{i}(u)\,\widehat{W}_{s}^{i}(z)^{\top}.$
Then
$\mathbb{E}\!\Big[W_{u}\!\int_{0}^{s}C(s,z)\,dW_{z}^{\top}\;\Big|\;\mathcal{F}_{s}^{i}\Big]=\int_{0}^{s}\partial_{z}\mathcal{M}_{s}^{i}(u,z)\,C(s,z)^{\top}\,dz.$
Splitting $\mathcal{M}_{s}^{i}=\mathsf{C}_{s}^{i}+\widehat{W}_{s}^{i}\,\widehat{W}_{s}^{i,\top}$ separates this into a covariance piece and a product-of-means piece:
$\int_{0}^{s}\partial_{z}\mathcal{M}_{s}^{i}(u,z)\,C(s,z)^{\top}\,dz=\int_{0}^{s}\partial_{z}\mathsf{C}_{s}^{i}(u,z)\,C(s,z)^{\top}\,dz\;+\;\widehat{W}_{s}^{i}(u)\!\int_{0}^{s}C(s,z)\,d\widehat{W}_{s}^{i}(z)^{\top}.$ (A.7)
Differentiating $\mathbb{E}[W_{u}\,\varphi_{t}^{r}]$ and collecting $dt$ terms then gives
$\frac{d}{dt}\mathbb{E}\!\big[W_{u}\,\varphi_{t}^{r}\big]$ $=i\,\mathbb{E}\!\bigg[\varphi_{t}^{r}\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,C(t,s)^{\top}\,ds\;+\;\widehat{W}_{t}^{i}(u)\!\int_{0}^{t}C(t,s)\,d\widehat{W}_{t}^{i}(s)^{\top}\;+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\Big)\bigg]\vec{r}(t).$ (A.8)
Coefficient matching. Equating (A.6) and (A.8) for all $\vec{r}(\cdot)$ and all $t$ yields two conditions. Matching the terms multiplying $\vec{r}(t)$ (after canceling the common $\widehat{W}_{t}^{i}(u)\int C\,d\widehat{W}^{i,\top}$ terms that appear on both sides via (A.7)) gives (A.3). Requiring the residual drift in (A.6) to vanish gives
$\int_{0}^{t}\big(B(u;t,s)+B^{Y}(u,t)\,C(t,s)\big)\,d\widehat{W}_{t}^{i}(s)=0,$
hence $B(u;t,s)=-B^{Y}(u,t)\,C(t,s)$, which is (A.4). ∎
A.2 Specialization to the physical observation model
We now specialize to the observation equation of Section 2,
$dY_{t}^{i}=\sqrt{P^{i}(t)}\,X_{t}\,dt+E^{i}\,dW_{t},$
which corresponds to setting $C(t,s)=\sqrt{P^{i}(t)}\,X_{t}(s)$ in the general framework of Proposition A.2.
Innovation process.
Define the innovation
$dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt,$
where $\widehat{X}_{t}^{i}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{i}]$. By the innovations theorem for Gaussian linear filtering (Liptser–Shiryaev [16], Theorem 7.12), $I^{i}$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion.
Innovation form of the dynamics.
Substituting $C(t,s)=\sqrt{P^{i}(t)}\,X_{t}(s)$ into Proposition A.2 and using $dI_{t}^{i}=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$, the dynamics (A.2) become
$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,X_{t}(s)^{\top}\,ds\Big)\sqrt{P^{i}(t)}\,dI_{t}^{i}\,+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\,dI_{t}^{i}.$ (A.9)
Identifying $\widetilde{X}^{i}$ as the filtering gain.
For $u<t$, the indicator term in (A.9) is locally constant in $u$, so the mixed $(t,u)$-update takes the form
$d_{t}\!\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du,\qquad 0\leq u<t\leq T.$ (A.10)
To verify the gain coefficient, recall that $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$ and that the innovation (A.17) has martingale component $E^{i}\,dW_{t}$ and absolutely continuous component $\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}\,dt$. The gain in (A.10) equals the conditional cross-covariance density $\operatorname{Cov}(dW_{u},\,\sqrt{P^{i}(t)}(X_{t}{-}\widehat{X}_{t}^{i})\,dt\mid\mathcal{F}_{t}^{i})/dt$; by Itô isometry applied to $\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$, this picks out $\widetilde{X}^{i}_{t}(u)$. Thus $\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}$ is the Kalman gain for updating the $u$-th noise increment: it maps observation innovations at time $t$ into revisions of the noise estimate at the earlier time $u$.
Remark A.1 (Conditional cross-covariance interpretation).
The identification above can be restated as
$\widetilde{X}^{i}_{t}(u)\;=\;\frac{\operatorname{Cov}\!\big(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i}\big)}{du}.$
Integrating in $u$ recovers the full conditional cross-covariance
$\operatorname{Cov}\!\big(X_{t},\,W_{u}\mid\mathcal{F}_{t}^{i}\big)\;=\;\int_{0}^{u}\widetilde{X}^{i}_{t}(s)\,ds\;=\;\int_{0}^{t}X_{t}(s)\,\partial_{s}\mathsf{C}_{t}^{i}(s,u)\,ds,$ (A.11)
where the second equality follows from $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}X_{t}(s)\,d\widetilde{W}_{t}^{i}(s)$ and Itô isometry against $\widetilde{W}_{t}^{i}(u)$.
Algebraic identity for $\widetilde{X}^{i}$ in terms of $F^{i}$.
Recall the induced estimated-state kernel $\widehat{X}^{i}_{t}(s):=X_{t}(s)\Pi^{i}+\int_{0}^{t}X_{t}(z)\,F_{t}^{i}(z,s)\,dz$. Since $\widetilde{X}^{i}_{t}(u):=X_{t}(u)-\widehat{X}^{i}_{t}(u)$,
$\widetilde{X}^{i}_{t}(u)=X_{t}(u)(I-\Pi^{i})-\int_{0}^{t}X_{t}(z)\,F_{t}^{i}(z,u)\,dz.$ (A.12)
This identity is algebraic: once $F^{i}$ is determined (Section A.3), $\widetilde{X}^{i}$ follows with no further fixed-point step.
Remark A.2 (State uncertainty as a contraction).
The conditional state covariance satisfies
$\Sigma^{XX,i}(t)=\int_{0}^{t}X_{t}(s)\,\widetilde{X}^{i}_{t}(s)^{\top}\,ds,$ (A.13)
which follows from Itô isometry applied to $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$.
Remark A.3 (Kalman-filter interpretation).
Identity (A.13) says that the $d\times d$ state uncertainty $\Sigma^{XX,i}(t)$ is obtained by “projecting” the unresolved kernel $\widetilde{X}^{i}_{t}(s)$ back through the state impulse response $X_{t}(s)$. In the single-player case with no endogenous drift, this reduces to the classical relation between the Kalman gain and the error covariance. In the multi-player setting the identity persists because the forward kernel environment is deterministic.
A.3 The filtering kernel $F^{i}$
We define $F^{i}$ as the primitive filtering object through its evolution equation. The unresolved kernel $\widetilde{X}^{i}$ is then a derived quantity via the algebraic identity (A.12).
Definition A.1 (Filtering kernel system).
Fix player $i$. A deterministic causal kernel $F^{i}$ on $\Delta_{T}$ solves the filtering kernel system if, with $\widetilde{X}^{i}$ defined by (A.12), it satisfies $F_{0}^{i}=0$ and
$\partial_{t}F_{t}^{i}(u,s)$ $=\widetilde{X}^{i}_{t}(u)^{\top}P^{i}(t)\,\widetilde{X}^{i}_{t}(s),\qquad u,s<t,$ (A.14)
$F_{t}^{i}(u,t)$ $=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}_{t}(u).$ (A.15)
Since (A.12) expresses $\widetilde{X}^{i}$ as a deterministic linear functional of $F^{i}$, the system (A.14)–(A.15) is a closed forward ODE for $F^{i}$ (quadratic through the substitution).
Proposition A.3 (Explicit formula).
Integrating (A.14) from $\max(u,s)$ to $t$ and adding the boundary contributions gives
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.16)
The three terms are: (i) the unresolved response at observation time $s$, projected through the observation channel; (ii) the drift component of the innovation at time $u$, revealing information about earlier increments at $s\leq u$; (iii) accumulated indirect inference from observations at times $r\in[\max(u,s),\,t]$.
Remark A.4 (Self-adjointness).
Under $(u,s)\mapsto(s,u)$ with transpose, the first two terms in (A.16) exchange and the integral is invariant (since $P^{i}(r)$ is symmetric). Hence $F_{t}^{i}(u,s)=F_{t}^{i}(s,u)^{\top}$, confirming that the induced map is an orthogonal projection.
Remark A.5 (Uniqueness).
If $F^{i}$ and $\tilde{F}^{i}$ both yield the density representation of Theorem 3.1(i), Itô isometry gives $F_{t}^{i}(u,s)=\tilde{F}_{t}^{i}(u,s)$ a.e.
Remark A.6 (Computation).
In the numerical scheme of Section 4, $\widetilde{X}^{i}$ is computed from $F^{i}$ via (A.12) at each time step, and the integral in (A.16) is accumulated forward alongside the $F^{i}$ update.
We now convert the innovations-based characterization of Subsections A.1A.2 into primitive-noise coordinates. The goal is to construct the deterministic kernel $F^{i}$ such that for every $t\in[0,T]$ and $u<t$,
$d_{u}\widehat{W}_{t}^{i}(u)=\Pi^{i}\,dW_{u}+\Bigg(\int_{0}^{t}F_{t}^{i}(u,s)\,dW_{s}\Bigg)\,du.$
A.3.1 Derivation
Recall the innovation form of the estimated-noise dynamics (equation (A.9)): for $u\leq t$,
$d_{u}\widehat{W}_{t}^{i}(u)=E^{i,\top}\,dI_{u}^{i}+\Bigg(\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}\Bigg)\,du,$
where the first term is the direct observation at time $u$ and the second is the accumulated drift-based revision over $(u,t]$. To pass from innovations to primitive noise, we expand each $dI_{s}^{i}$ in terms of $dW$.
Expanding the innovation.
The physical-measure decomposition of the innovation is
$dI_{t}^{i}=E^{i}\,dW_{t}\;+\;\sqrt{P^{i}(t)}\,(X_{t}-\widehat{X}_{t}^{i})\,dt.$
The first piece is the martingale component; the second is absolutely continuous. Substituting the linear forms and using $\widetilde{X}^{i}_{t}(r):=X_{t}(r)-\widehat{X}^{i}_{t}(r)$ (the definition of $\widetilde{X}^{i}$ and (A.13)):
$dI_{t}^{i}=E^{i}\,dW_{t}+\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}_{t}(r)\,dW_{r}\,dt.$ (A.17)
Expanding the direct term.
Substituting (A.17) at $t=u$ into $E^{i,\top}\,dI_{u}^{i}$:
$E^{i,\top}\,dI_{u}^{i}$ $=\underbrace{E^{i,\top}E^{i}\,dW_{u}}_{=\;\Pi^{i}\,dW_{u}}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\int_{0}^{u}\widetilde{X}^{i}_{u}(r)\,dW_{r}\,du.$ (A.18)
The first piece contributes the $\Pi^{i}\,dW_{u}$ in the density representation. The second is absolutely continuous in $u$ with primitive-noise density $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)$ at source-noise index $s\leq u$.
Expanding the integral term.
Substituting (A.17) into $\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$:
$\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$ $=\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\,dW_{s}}_{\text{(a) martingale}}$
$\quad+\;\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}P^{i}(s)\int_{0}^{s}\widetilde{X}^{i}_{s}(r)\,dW_{r}\,ds}_{\text{(b) abs.\ continuous}}.$
Term (a) contributes $\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$ at source-noise index $s\in(u,t]$. Applying stochastic Fubini to term (b) and reading off the coefficient of $dW_{r}$:
$\text{(b)}=\int_{0}^{t}\Bigg(\int_{\max(u,r)}^{t}\widetilde{X}^{i}_{s}(u)^{\top}P^{i}(s)\,\widetilde{X}^{i}_{s}(r)\,ds\Bigg)\,dW_{r}.$
Assembling the density.
Collecting the coefficient of $dW_{s}\,du$ from all three contributions and noting that causality ($\widetilde{X}^{i}_{s}(u)=0$ for $u>s$) automatically zeroes the irrelevant indicators:
$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.19)
The three terms admit a direct interpretation:
  • $\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$: the unresolved response at observation time $s$, projected through the observation channel. Nonzero only when $u\leq s$ (causal in the observation direction).
  • $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)$: the drift component of the innovation at time $u$, which reveals information about earlier noise increments at index $s\leq u$. Nonzero only when $s\leq u$ (causal in the noise direction).
  • $\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr$: accumulated indirect inference—the joint information about the $u$-th and $s$-th increments extracted from observations at times $r\in[\max(u,s),\,t]$.
Remark A.7 (Self-adjointness and uniqueness).
Under $(u,s)\mapsto(s,u)$ with transpose, the first two terms exchange and the integral is invariant (since $P^{i}(r)$ is symmetric). Hence $F_{t}^{i}(u,s)=F_{t}^{i}(s,u)^{\top}$, confirming that the induced map $(\mathcal{P}_{t}^{i}f)(s):=\Pi^{i}f(s)+\int_{0}^{t}F_{t}^{i}(r,s)^{\top}f(r)\,dr$ is self-adjoint, as required by its interpretation as an orthogonal projection. Uniqueness follows from Itô isometry: if $\tilde{F}$ also satisfies the density representation, then $\int_{0}^{t}\|F_{t}^{i}(u,s)-\tilde{F}_{t}^{i}(u,s)\|^{2}\,ds=0$ a.e.
Remark A.8 (Boundary condition and evolution).
Evaluating (A.19) at $s=t$ (using $\widetilde{X}^{i}_{u}(t)=0$ for $u<t$):
$F_{t}^{i}(u,t)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}_{t}(u),$
which are each other’s transposes. Differentiating (A.19) in $t$ at fixed $u,s<t$ gives the evolution
$\partial_{t}F_{t}^{i}(u,s)=\widetilde{X}^{i}_{t}(u)^{\top}P^{i}(t)\,\widetilde{X}^{i}_{t}(s),$
which is manifestly self-adjoint under $(u,s)\mapsto(s,u)^{\top}$.
Remark A.9 (Computational simplification).
The formula (A.19) is explicit in the gain kernel $\widetilde{X}^{i}$: once $\widetilde{X}^{i}$ is computed from (A.12), the filtering kernel $F^{i}$ requires no further fixed-point iteration. In the numerical scheme of Section 4, the integral $\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr$ is accumulated forward alongside the $\widetilde{X}^{i}$-computation, and the two boundary terms are set at each new observation time.
A.3.2 Formal characterization
Definition A.2 (Filtering kernel system (definition of $F^{i}$)).
Fix player $i$. The inputs are the state kernel $X(\cdot,\cdot)$ and the unresolved kernel $\widetilde{X}^{i}$ from Subsection A.2. The output is the primitive-noise density kernel $F^{i}$ of Theorem 3.1(i).
A deterministic kernel $F^{i}$ is said to be admissible if it is causal ($F_{t}^{i}(u,s)=0$ when $\max\{u,s\}>t$) and, together with the induced kernel
$\begin{split}F_{t}^{i}(u,s)&=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\\ &\quad+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr,\qquad u,s\leq t.\end{split}$ (A.20)
We define $F^{i}$ to be the unique causal solution of (A.12)–(A.20).
Proposition A.4 (Verification: primitive-noise density of the noise-state).
Let $F^{i}$ be the kernel defined by Definition A.1. Then for every $t\in[0,T]$ and $u<t$,
$d_{u}\widehat{W}_{t}^{i}(u)=\Pi^{i}\,dW_{u}+\Bigg(\int_{0}^{t}F_{t}^{i}(u,s)\,dW_{s}\Bigg)\,du.$
Moreover, $F^{i}$ is unique among deterministic causal kernels with this property.
Proof of uniqueness.
If $F^{i}$ and $\tilde{F}^{i}$ both satisfy the density representation, then $\int_{0}^{t}(F_{t}^{i}(u,s)-\tilde{F}_{t}^{i}(u,s))\,dW_{s}=0$ a.s. for each $(t,u)$. By Itô isometry, $\int_{0}^{t}\|F_{t}^{i}(u,s)-\tilde{F}_{t}^{i}(u,s)\|^{2}\,ds=0$, hence $F_{t}^{i}(u,s)=\tilde{F}_{t}^{i}(u,s)$ a.e. ∎
Appendix B Control First Order Conditions
This appendix derives the stationarity condition (B.8) and the closed backward system for the adjoint kernels $(\bar{H}^{X},H^{X},\{\bar{H}^{k},H^{k}\}_{k\neq i})$. By the innovations exogeneity established in Corollary 3.5, the best-response problem for player $i$ (opponents’ maps fixed) is an optimization over deterministic impulse-response maps weighting the exogenous innovation $I^{i}$. [^2] To avoid circularity: opponents’ deterministic kernels make $I^{i}$ a standard Brownian motion independent of $D^{i}$ [16, Theorem 7.12]; Lemma 3.4 then gives strict convexity over the full $L^{2}$ class; the FOC derived here is therefore necessary and sufficient, yielding the deterministic-kernel best response of Corollary 3.5.
B.1 The best-response problem in innovations coordinates
Fix a baseline noise-state linear profile and a deviating player $i$. By the innovations exogeneity established in Corollary 3.5, $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$, and $I^{i}$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion invariant to $D^{i}$. Player $i$’s control
$D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$
is a deterministic affine functional of $I^{i}$, and the optimization is over $(\bar{D}^{i},D^{i})$ with no self-referential signal dependence. Since the state dynamics are linear in $D^{i}$ and the cost is quadratic with $G^{DD,i}(t)\succ 0$, the induced objective is strictly convex: the stationarity condition derived below is the unique global optimum.
B.2 Spike variation and the first-variation system
We use spike variations to expose the adjoint structure. Fix $t\in[0,T)$ and $v\in L^{2}(\Omega,\mathcal{F}_{t}^{i};\mathbb{R}^{d})$. The spike perturbation $D_{s}^{i,\rho,\varepsilon}:=D_{s}^{i}+\rho\,v\,\mathbf{1}_{[t,t+\varepsilon]}(s)$ induces the normalized first variation $\delta X_{s}:=\lim_{\varepsilon\downarrow 0}\frac{1}{\varepsilon}\frac{\partial}{\partial\rho}X_{s}^{\rho,\varepsilon}\big|_{\rho=0}$ with $\delta X_{t}=v$.
Because opponents apply fixed linear maps with deterministic kernels, the map $v\mapsto(\delta X_{s},\{\delta w_{s}^{k}(\cdot)\}_{k\neq i})$ is exactly linear in $v$ with deterministic coefficients—not a linearization but the full response. The deviator’s own innovation is unaffected ($\delta I_{s}^{i}=0$), since $\delta X_{s}\in\mathcal{F}_{t}^{i}$ implies $\delta\widehat{X}_{s}^{i}=\delta X_{s}$.
Opponents’ filter response.
For each $k\neq i$, $\delta(d_{u}\widehat{W}_{s}^{k}(u))=\delta w_{s}^{k}(u)\,du$ with $\delta w_{t}^{k}(\cdot)\equiv 0$ and
$\partial_{s}\delta w_{s}^{k}(u)=\widetilde{X}^{k}_{s}(u)^{\top}P^{k}(s)\Bigl(\delta X_{s}-\int_{0}^{s}X_{s}(r)\,\delta w_{s}^{k}(r)\,dr\Bigr),\qquad 0\leq u\leq s.$ (B.1)
The opponent control variation is $\delta D_{s}^{k}=\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du$, giving the coupled state system
$\frac{d}{ds}\delta X_{s}=A(s)\,\delta X_{s}+\sum_{k\neq i}\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du,\qquad\delta X_{t}=v.$ (B.2)
B.3 Transition kernels
The coupled system (B.1)–(B.2) has bounded deterministic coefficients and admits a unique deterministic two-parameter evolution family.
Definition B.1 (Block transition kernels).
For $s\geq t$, define $\Phi^{XX}(s,t)\in\mathbb{R}^{d\times d}$ and $\Phi^{Xk}(s,t,u)$ ($k\neq i$, $u\in[0,t]$) by
$\delta X_{s}=\Phi^{XX}(s,t)\,v+\sum_{k\neq i}\int_{0}^{t}\Phi^{Xk}(s,t,u)\,\eta^{k}(u)\,du,$
for initial data $\delta X_{t}=v$, $\delta w_{t}^{k}(\cdot)=\eta^{k}(\cdot)$. In the spike deviation $\eta^{k}\equiv 0$, so $\delta X_{s}=\Phi^{XX}(s,t)\,v$.
These blocks satisfy
$\frac{\partial}{\partial t}\Phi^{XX}(s,t)$ $=-\Phi^{XX}(s,t)A(t)-\sum_{k\neq i}\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}_{t}(z)^{\top}P^{k}(t)\,dz,$ (B.3)
$\frac{\partial}{\partial t}\Phi^{Xk}(s,t,u)$ $=-\Phi^{XX}(s,t)D^{k}_{t}(u)+\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}_{t}(z)^{\top}P^{k}(t)\,X_{t}(u)\,dz,$ (B.4)
with $\Phi^{XX}(t,t)=I$ and $\Phi^{Xk}(t,t,u)=0$.
B.4 First variation of costs and the adjoint kernels
Fix player $i$ and suppress player superscripts on weight matrices. Using $\delta X_{s}=\Phi^{XX}(s,t)v$ and the linear form of $X_{s}$, the normalized cost variation is
$\delta J_{t}^{i}(v)=2\,\mathbb{E}\!\left[\left.v^{\top}\Bigl(G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X}_{t}+\int_{0}^{T}H^{X}_{t}(u)\,d\widehat{W}_{t}^{i}(u)\Bigr)\ \right|\mathcal{F}_{t}^{i}\right],$ (B.5)
where
$\bar{H}^{X}_{t}$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}\bigl[G^{XX}(s)\bar{X}_{s}+\bar{G}^{X,i}_{s}\bigr]\,ds+\Phi^{XX}(T,t)^{\top}\bigl[G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}\bigr],$ (B.6)
$H^{X}_{t}(u)$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}\bigl[G^{XX}(s)\,X_{s}(u)+G^{X,i}_{s}(u)\bigr]\,ds+\Phi^{XX}(T,t)^{\top}\bigl[G^{XX,i}(T)\,X_{T}(u)+G^{X,i}_{T}(u)\bigr].$ (B.7)
Stationarity condition (globally necessary and sufficient).
Setting $\delta J_{t}^{i}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives
$G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X}_{t}+\int_{0}^{T}H^{X}_{t}(u)\,d\widehat{W}_{t}^{i}(u)=0\qquad\text{a.s.,}$ (B.8)
with equilibrium coefficients $\bar{D}^{i}_{t}=-G^{DD,i}(t)^{-1}\bar{H}^{X}_{t}$ and $D^{i}_{t}(u)=-G^{DD,i}(t)^{-1}H^{X}_{t}(u)$.
Equivalently, $D_{t}^{i}=-(G^{DD,i}(t))^{-1}\mathbb{E}[\bar{H}^{X}_{t}+\int_{0}^{T}H^{X}_{t}(u)\,dW_{u}\mid\mathcal{F}_{t}^{i}]$: the control is the conditional expectation of the full-information action (a certainty-equivalence property).
B.5 Closed backward system for the adjoint kernels
Define the belief-adjoint objects
$\bar{H}^{k}_{t}(u)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}\bigl[G^{XX}(s)\bar{X}_{s}+\bar{G}^{X,i}_{s}\bigr]\,ds+\Phi^{Xk}(T,t,u)^{\top}\bigl[G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}\bigr],$ (B.9)
$H^{k}_{t}(u,r)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}\bigl[G^{XX}(s)\,X_{s}(r)+G^{X,i}_{s}(r)\bigr]\,ds$
$\qquad+\;\Phi^{Xk}(T,t,u)^{\top}\bigl[G^{XX,i}(T)\,X_{T}(r)+G^{X,i}_{T}(r)\bigr].$ (B.10)
Then $(\bar{H}^{X},\{\bar{H}^{k}\},H^{X},\{H^{k}\})$ satisfy
$\frac{d}{dt}\bar{H}^{X}_{t}$ $=-\bigl[G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\bigr]-A(t)^{\top}\bar{H}^{X}_{t}-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k}_{t}(z)\,dz,$ (B.11)
$\frac{d}{dt}\bar{H}^{k}_{t}(u)$ $=-D^{k}_{t}(u)^{\top}\bar{H}^{X}_{t}+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k}_{t}(z)\,dz,$ (B.12)
$\frac{d}{dt}H^{X}_{t}(r)$ $=-\bigl[G^{XX,i}(t)X_{t}(r)+G^{X,i}_{t}(r)\bigr]-A(t)^{\top}H^{X}_{t}(r)-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,H^{k}_{t}(z,r)\,dz,$ (B.13)
$\frac{d}{dt}H^{k}_{t}(u,r)$ $=-D^{k}_{t}(u)^{\top}H^{X}_{t}(r)+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,H^{k}_{t}(z,r)\,dz,$ (B.14)
with terminal conditions $\bar{H}^{X}_{T}=G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}$,  $\bar{H}^{k}_{T}(\cdot)=0$,  $H^{X}_{T}(r)=G^{XX,i}(T)X_{T}(r)+G^{X,i}_{T}(r)$,  $H^{k}_{T}(\cdot,r)=0$.
B.6 Risk-sensitive extension: exponential utility FOC
Theorem B.1 (Risk-sensitive extension).
Replace (2.4) with the entropic risk measure
$J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}\!\left[e^{\,\theta_{i}\mathcal{C}^{i}}\right],\qquad\theta_{i}>0,$ (B.15)
where $\mathcal{C}^{i}$ is the random cost (2.3). A Taylor expansion gives $J^{i}_{\theta_{i}}=\mathbb{E}[\mathcal{C}^{i}]+\tfrac{\theta_{i}}{2}\operatorname{Var}(\mathcal{C}^{i})+O(\theta_{i}^{2})$, so $\theta_{i}$ governs the degree of risk aversion (the risk-neutral case $\theta_{i}\to 0$ recovers (2.4)).
Under a linear profile, the best-response FOC is identical to Corollary 3.5 with $\widehat{W}_{t}^{i}$ replaced by the risk-adjusted noise-state
$\widehat{W}^{\,i,\theta}_{t}:=\bigl(I-\theta_{i}C_{t}K_{t}\bigr)^{-1}\!\left(\widehat{W}_{t}^{i}+\theta_{i}C_{t}k_{t}\right),$
where $C_{t}:=\mathsf{C}_{t}^{i}(u,s)=\operatorname{Cov}(W_{u},W_{s}\mid\mathcal{F}_{t}^{i})$ is the deterministic conditional covariance kernel of $W$ given $\mathcal{F}_{t}^{i}$ (deterministic by Gaussianity of the conditional law under the linear profile), $K_{t}$ is the symmetric quadratic kernel of $\mathcal{C}^{i}$ in the Wiener-chaos representation (B.17), and $k_{t}$ is the corresponding linear kernel (B.16). The formula is valid when $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$, which holds for $\theta_{i}$ sufficiently small or $T$ sufficiently short. The filtering structure is unchanged; certainty equivalence fails because the quadratic cost kernel rescales conditional precision via $(I-\theta_{i}C_{t}K_{t})^{-1}$ while $k_{t}$ shifts the conditional mean.
We now derive this result. Replace (2.4) with $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$, where $\mathcal{C}^{i}$ is the random cost (2.3). The spike variation of Section B.2 carries over; differentiating $J^{i}_{\theta_{i}}$ in the spike amplitude gives
$\delta J^{i}_{\theta_{i},t}(v)=\mathbb{E}\!\left[Z^{i}\,\delta\mathcal{C}^{i}_{t}(v)\right],\qquad Z^{i}:=\frac{e^{\,\theta_{i}\mathcal{C}^{i}}}{\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]},$
with $\delta\mathcal{C}^{i}_{t}(v)$ identical to (B.5).
Quadratic structure of the cost.
Expanding $\mathcal{C}^{i}$ via the linear forms $X_{s}=\bar{X}_{s}+\int_{0}^{s}X_{s}(u)\,dW_{u}$ and $D^{i}_{s}=\bar{D}^{i}_{s}+\int_{0}^{s}\mathcal{D}^{i}_{s}(u)\,dW_{u}$ and applying stochastic Fubini yields the Wiener-chaos representation
$\mathcal{C}^{i}=c_{0}+\int_{0}^{T}\ell(r)^{\top}dW_{r}+\tfrac{1}{2}\int_{0}^{T}\!\!\int_{0}^{T}dW_{r}^{\top}K(r,s)\,dW_{s},$
where $c_{0}$ absorbs deterministic and trace terms, and the linear and symmetric quadratic kernels are
$\ell(r)$ $:=\int_{r}^{T}\!\Big[X_{t}(r)^{\top}\bigl(G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\bigr)+G^{X,i}_{t}(r)^{\top}\bar{X}_{t}$
$\qquad\qquad+\mathcal{D}^{i}_{t}(r)^{\top}G^{DD,i}(t)\,\bar{D}^{i}_{t}\Big]\,dt$
$\qquad+\;X_{T}(r)^{\top}\bigl(G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}\bigr)+G^{X,i}_{T}(r)^{\top}\bar{X}_{T},$ (B.16)
$K(r,s)$ $:=\int_{\max(r,s)}^{T}\!\Big[X_{t}(r)^{\top}G^{XX,i}(t)\,X_{t}(s)+X_{t}(r)^{\top}G^{X,i}_{t}(s)$
$\qquad\qquad+G^{X,i}_{t}(r)^{\top}X_{t}(s)+\mathcal{D}^{i}_{t}(r)^{\top}G^{DD,i}(t)\,\mathcal{D}^{i}_{t}(s)\Big]\,dt$
$\qquad+\;X_{T}(r)^{\top}G^{XX,i}(T)\,X_{T}(s)+X_{T}(r)^{\top}G^{X,i}_{T}(s)+G^{X,i}_{T}(r)^{\top}X_{T}(s).$ (B.17)
Tilted conditional FOC.
Since $\mathcal{C}^{i}$ is quadratic in $W$, the weight $Z^{i}$ tilts a Gaussian measure by a quadratic form, preserving Gaussianity. Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ and applying the tower property gives
$G^{DD,i}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}\!\left[\mathcal{M}^{i}_{t}\,\Big|\,\mathcal{F}_{t}^{i}\right]=0,\qquad\mathbb{E}^{Q^{i}}[\,\cdot\mid\mathcal{F}_{t}^{i}]:=\frac{\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]}{\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]},$
which has the structure of (B.8) under the tilted measure.
Risk-adjusted noise-state.
Since $\mathcal{M}^{i}_{t}$ is linear in $W$ with deterministic kernels, the FOC reduces to computing the tilted conditional mean $\widehat{W}^{\,i,\theta}_{t}:=\mathbb{E}^{Q^{i}}[W_{r}\mid\mathcal{F}_{t}^{i}]$. Under the deterministic-kernel profile, $\mathcal{L}(W\mid\mathcal{F}_{t}^{i})$ is Gaussian with mean $m_{t}:=\widehat{W}_{t}^{i}$ and deterministic covariance $C_{t}:=\mathsf{C}_{t}^{i}$. The quadratic tilt shifts this to
$C_{t,\theta}^{-1}=C_{t}^{-1}-\theta_{i}K_{t},\qquad m_{t,\theta}=C_{t,\theta}(C_{t}^{-1}m_{t}+\theta_{i}k_{t}),$
provided $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$ (satisfied for $|\theta_{i}|$ small or $T$ short). Here $C_{t}$ and $K_{t}$ act as integral operators on $L^{2}([0,t];\mathbb{R}^{(n+1)d})$, so $(I-\theta_{i}C_{t}K_{t})^{-1}$ is the resolvent (Neumann series) $(I-\theta_{i}C_{t}K_{t})^{-1}f=f+\theta_{i}C_{t}K_{t}f+\theta_{i}^{2}(C_{t}K_{t})^{2}f+\cdots$, convergent precisely when $\|\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\|_{\mathrm{op}}<1$. Rearranging gives
$\widehat{W}^{\,i,\theta}_{t}=(I-\theta_{i}C_{t}K_{t})^{-1}(\widehat{W}_{t}^{i}+\theta_{i}C_{t}k_{t}),$
and substituting into the tilted FOC yields
$\boxed{G^{DD,i}(t)D^{i}_{t}\;+\;\bar{H}^{X}_{t}\;+\;\int_{0}^{T}H^{X}_{t}(r)\,d_{r}\,\widehat{W}^{\,i,\theta}_{t}(r)\;=\;0,}$ (B.18)
where
$\widehat{W}^{\,i,\theta}_{t}(r):=\bigl[(I-\theta_{i}C_{t}K_{t})^{-1}\bigl(\widehat{W}_{t}^{i}+\theta_{i}\,C_{t}k_{t}\bigr)\bigr](r).$ (B.19)
The risk-adjusted noise-state $\widehat{W}^{\,i,\theta}_{t}$ differs from the risk-neutral noise-state $\widehat{W}_{t}^{i}$ in two ways: the linear kernel $k_{t}$ (B.16) shifts the conditional mean of $W$ in the direction that reduces expected cost, while the quadratic kernel $K_{t}$ (B.17) rescales conditional precision through $(I-\theta_{i}C_{t}K_{t})^{-1}$, amplifying the weight on directions of $W$-space along which cost variance is large. As $\theta_{i}\to 0$, $\widehat{W}^{\,i,\theta}_{t}\to\widehat{W}_{t}^{i}$ and (B.18) recovers (B.8).
The first variation $\delta J^{i}_{\theta_{i},t}(v)$ is computed by differentiating $J^{i}_{\theta_{i}}$ in the spike amplitude (Section B.2), which introduces the exponential tilt $Z^{i}:=e^{\theta_{i}\mathcal{C}^{i}}/\mathbb{E}[e^{\theta_{i}\mathcal{C}^{i}}]$. Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives the tilted stationarity condition $G^{DD,i}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}[M^{i}_{t}\mid\mathcal{F}_{t}^{i}]=0$, where $\mathbb{E}^{Q^{i}}[\cdot\mid\mathcal{F}_{t}^{i}]:=\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]/\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]$ is the conditional expectation under the cost-tilted measure. Since $\mathcal{C}^{i}$ is quadratic in $W$ under the linear profile, $Q^{i}$ is a Gaussian measure, and computing the tilted conditional mean of the stochastic costate $M^{i}_{t}$ yields (B.18).
B.7 Information cost: proof of Proposition 3.6
Full- and partial-information controls.
At the equilibrium profile $D^{*}$, the stochastic costate is
$M^{i}_{t}:=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}W_{u},$
where $(\bar{H}^{X,i},H^{X,i})$ solve the backward system (B.11)–(B.14) evaluated at $D^{*}$. The full-information and equilibrium controls are
$D^{i,\mathrm{full}}_{t}$ $=-(G^{DD,i}(t))^{-1}M^{i}_{t},$ (B.20)
$D^{i}_{t}$ $=-(G^{DD,i}(t))^{-1}\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}].$ (B.21)
Both are evaluated at the same adjoint kernels: opponents’ maps are fixed and linear with deterministic kernels, so the backward system is identical regardless of player $i$’s information. The control gap is
$\delta D^{i}_{t}:=D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$
where $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$.
Completion of the square.
Since $D^{i,\mathrm{full}}$ minimizes the strictly convex quadratic $D^{i}\mapsto J^{i}(D^{i};D^{*,-i})$ over $L^{2}(\mathcal{F})$ (Lemma 3.4), and $\delta D^{i}\in L^{2}(\mathcal{F})$, the first variation vanishes and
$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}(\delta D^{i}_{t})^{\top}G^{DD,i}(t)\,\delta D^{i}_{t}\,\mathrm{d}t\right].$ (B.22)
To verify that only $G^{DD,i}$ appears: since $\delta X$ is affine in $\delta D^{i}$ (the dynamics are linear), the full cost perturbation contains terms of order zero, one, and two in $\delta D^{i}$. The zero-order term vanishes by definition. All first-order terms—including those from $\delta X^{\top}G^{i}_{XX}X^{\mathrm{full}}$ and $(G^{X,i})^{\top}\delta X$—vanish by the stationarity condition at $D^{i,\mathrm{full}}$. The second-order terms comprise the direct control cost $(\delta D^{i})^{\top}G^{DD,i}\,\delta D^{i}$ and the state cost $\delta X^{\top}G^{i}_{XX}\,\delta X$ (plus terminal). The standard LQG completion-of-the-square identifies these: using $G^{DD,i}(t)\,D^{i,\mathrm{full}}_{t}+M^{i}_{t}=0$ and the definition of $M^{i}_{t}$ through the transition kernel $\Phi_{XX}$ (which incorporates opponents’ linear feedback), the state-cost second-order terms cancel against cross-terms from the adjoint, leaving (B.22).
Kernel formula for $\Sigma^{M,i}_{t}$.
Substituting the control gap into (B.22):
$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}\left(\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u)\right)^{\top}(G^{DD,i}(t))^{-1}\left(\int_{0}^{t}H^{X,i}_{t}(v)\,\mathrm{d}\widetilde{W}^{i}_{t}(v)\right)\mathrm{d}t\right].$ (B.23)
Since $\widetilde{W}^{i}_{t}$ is Gaussian with deterministic covariance $C^{i}_{t}(u,v)$, the expectation evaluates to $\mathrm{tr}[(G^{DD,i}(t))^{-1}\,\Sigma^{M,i}_{t}]$ at each $t$, giving (3.7). It remains to verify the kernel representation of $\Sigma^{M,i}_{t}$.
The conditional mean of $M^{i}_{t}$ given $\mathcal{F}^{i}_{t}$ is
$\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\bar{H}^{X,i}_{t}+\int_{0}^{t}\widehat{H}^{X,i}_{t}(s)\,\mathrm{d}W_{s},$
where the resolved adjoint kernel is
$\widehat{H}^{X,i}_{t}(s):=H^{X,i}_{t}(s)\,\Pi^{i}+\int_{s}^{t}H^{X,i}_{t}(u)\,F^{i}_{t}(u,s)\,\mathrm{d}u.$ (B.24)
This is obtained by substituting the noise-state decomposition $\mathrm{d}_{u}\widehat{W}^{i}_{t}(u)=\Pi^{i}\,\mathrm{d}W_{u}+\bigl(\int_{0}^{t}F^{i}_{t}(u,s)\,\mathrm{d}W_{s}\bigr)\,\mathrm{d}u$ (Theorem 3.1(i)) into $\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widehat{W}^{i}_{t}(u)$ and applying stochastic Fubini. Then
$M^{i}_{t}-\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\int_{0}^{t}\widetilde{H}^{X,i}_{t}(s)\,\mathrm{d}W_{s},$
where the unresolved adjoint kernel $\widetilde{H}^{X,i}_{t}(s):=H^{X,i}_{t}(s)-\widehat{H}^{X,i}_{t}(s)=H^{X,i}_{t}(s)(I-\Pi^{i})-\int_{s}^{t}H^{X,i}_{t}(u)\,F^{i}_{t}(u,s)\,\mathrm{d}u$ is the analogue of $\widetilde{X}^{i}$ for the backward system. By Itô isometry,
$\Sigma^{M,i}_{t}=\int_{0}^{t}\widetilde{H}^{X,i}_{t}(s)\,\widetilde{H}^{X,i}_{t}(s)^{\top}\,\mathrm{d}s.$
Separation from control.
The covariance $C^{i}_{t}(u,v)$ of $\widetilde{W}^{i}_{t}$ is deterministic (Gaussianity of the conditional law) and depends on the signal structure $(P^{i},X)$ but not on the realized signal path or on $D^{i}$. The adjoint $H^{X,i}$ depends on the equilibrium profile $D^{*,-i}$ through the forward environment, not on $D^{i}$ itself. Both $\widetilde{H}^{i}_{X}$ and $\Sigma^{M,i}_{t}$ are therefore deterministic functionals of $(P^{i},D^{*,-i})$, completing the proof. ∎
Appendix C Well-posedness of the deterministic impulse-response fixed point
This appendix formulates the equilibrium as a dynamic programming problem on the impulse-response state space, records the a priori bounds that follow from the projection structure of the filtering map, and establishes well-posedness: contraction for short horizons, existence for arbitrary horizons.
Notation.
$C$ denotes a generic constant depending on $(n,d,L,\lambda)$; $C(M)$ when it also depends on a policy-norm bound $M$.
C.1 Standing assumptions
Assumption C.1.
There exist $L>0$ and $\lambda>0$ such that, for all $t\in[0,T]$ and all $i$, $\|A(t)\|+\|\Sigma(t)\|+\|P^{i}(t)\|+\|G^{XX,i}(t)\|+\|\bar{G}^{X,i}_{t}\|+\|G^{X,i}(t,\cdot)\|_{L^{\infty}}+\|G^{XX,i}(T)\|+\|\bar{G}^{X,i}_{T}\|+\|G^{X,i}(T,\cdot)\|_{L^{\infty}}\leq L$ and $G^{DD,i}(t)\succeq\lambda\,I_{d}$.
C.2 The kernel state space and its dynamics
At time $t$ the kernel state is
$\mathcal{Z}_{t}:=\bigl(\bar{X}_{t},\;X_{t}(\cdot),\;\{F_{t}^{j}\}_{j=1}^{n}\bigr)\;\in\;\mathbb{R}^{d}\times L^{2}([0,t];\mathbb{R}^{d\times(n+1)d})\times\prod_{j=1}^{n}L^{2}_{\mathrm{sym}}([0,t]^{2};\mathbb{R}^{(n+1)d\times(n+1)d}).$ (C.1)
The unresolved kernel $\widetilde{X}^{j}$ is not a state variable; it is determined algebraically from $(\mathcal{Z}_{t})$ by
$\widetilde{X}^{j}_{t}(u)=X_{t}(u)(I-\Pi^{j})-\int_{0}^{t}X_{t}(z)\,F_{t}^{j}(z,u)\,dz.$ (C.2)
Under a noise-state linear profile $\mathbf{D}=\{(\bar{D}^{j},D^{j})\}_{j=1}^{n}$, the kernel state evolves by
$\dot{\bar{X}}(t)$ $=A(t)\bar{X}_{t}+\textstyle\sum_{k}\bar{D}^{k}_{t},$ (C.3)
$\dot{X}_{t}(s)$ $=A(t)X_{t}(s)+\textstyle\sum_{k}\mathcal{D}^{k}_{t}(s),\qquad X(s,s)=\Sigma(s),$ (C.4)
$\partial_{t}F_{t}^{j}(u,s)$ $=\widetilde{X}^{j}_{t}(u)^{\top}P^{j}(t)\,\widetilde{X}^{j}_{t}(s),\qquad F_{0}^{j}=0,$ (C.5)
with boundary $F_{t}^{j}(u,t)=\widetilde{X}^{j}_{t}(u)^{\top}\sqrt{P^{j}(t)}\,E^{j}$, and primitive-noise control $\mathcal{D}^{k}_{t}(s):=D^{k}_{t}(s)\Pi^{k}+\int_{0}^{t}D^{k}_{t}(u)\,F_{t}^{k}(u,s)\,du$.
The policy $\mathbf{D}$ enters (C.3)–(C.4) but does not enter (C.5) directly: $F^{j}$ depends on $\mathbf{D}$ only through $X$ via (C.2).
C.3 The Hamilton–Jacobi–Bellman equation
Let $V^{i}(t,\mathcal{Z})$ denote the equilibrium continuation cost for player $i$ from kernel state $\mathcal{Z}$ at time $t$. Since the kernel-state dynamics are deterministic, the HJB is a first-order terminal-value problem:
$\partial_{t}V^{i}+D_{\bar{X}}V^{i}\cdot\dot{\bar{X}}+\langle D_{X}V^{i},\,\dot{X}\rangle+\sum_{j=1}^{n}\langle D_{F^{j}}V^{i},\,\dot{F}^{j}\rangle+\ell^{i}(t,\mathcal{Z},\mathbf{D})=0,\qquad V^{i}(T,\cdot)=g^{i}(\cdot),$ (C.6)
where $\langle\cdot,\cdot\rangle$ denotes the $L^{2}$ pairing, $\ell^{i}$ is the running cost, and $g^{i}$ is the terminal cost, both evaluated at the kernel state.
Player $i$’s control $(\bar{D}^{i}_{t},D^{i}_{t}(\cdot))$ enters $\dot{\bar{X}}$ and $\dot{X}$ linearly but does not enter $\dot{F}^{j}$; the Hamiltonian minimization over $D^{i}$ therefore involves only $D_{\bar{X}}V^{i}$, $D_{X}V^{i}$, and the quadratic effort cost, giving the first-order condition
$\bar{D}^{i}_{t}=-(G^{DD,i}(t))^{-1}D_{\bar{X}}V^{i},\qquad D^{i}_{t}(u)=-(G^{DD,i}(t))^{-1}D_{X}V^{i}(u).$ (C.7)
Along the equilibrium path, $D_{\bar{X}}V^{i}=\bar{H}^{X,i}_{t}$ and $D_{X}V^{i}(\cdot)=H^{X,i}_{t}(\cdot)$, recovering Corollary 3.5.
The information wedge as a shadow price.
The $\sum_{j}\langle D_{F^{j}}V^{i},\dot{F}^{j}\rangle$ term in (C.6) captures the indirect effect of the kernel state’s evolution on player $i$’s continuation cost through opponents’ filters. Along the equilibrium path, differentiating through the chain $D^{i}\to X\to\widetilde{X}^{k}\to F^{k}$ identifies
$\bar{\mathcal{V}}^{i}_{t}=\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz=\sum_{k\neq i}\langle D_{F^{k}}V^{i},\,\partial_{X}\dot{F}^{k}\rangle,$ (C.8)
i.e. the mean information wedge of Theorem 3.2 is the shadow price of the opponent’s filter evolution, contracted against the sensitivity of that evolution to the state kernel. The kernel wedge $\mathcal{V}^{i}_{t}(r)$ admits the same interpretation with $\bar{H}^{k,i}$ replaced by $H^{k,i}(\cdot,\cdot,r)$.
Endogenous precision.
If $P^{i}(\cdot)$ is player $i$’s control subject to attention cost $\int_{0}^{T}c^{i}(t,P^{i}(t))\,dt$, then $P^{i}$ enters (C.6) only through $\dot{F}^{i}$ via (C.5), and the FOC
$\bigl\langle D_{F^{i}}V^{i},\;\widetilde{X}^{i}_{t}(\cdot)^{\top}(\cdot)\,\widetilde{X}^{i}_{t}(\cdot)\bigr\rangle=\frac{\partial c^{i}}{\partial P^{i}}(t,P^{i}(t))$ (C.9)
decouples from the FOC for $D^{i}$: the left side is the marginal value of accelerating the filter, which vanishes as $\mathcal{P}_{t}^{i}\to I$ (Lemma C.3), guaranteeing an interior solution when $\partial c^{i}/\partial P^{i}>0$.
C.4 A priori bounds from projection structure
The filtering kernel $F_{t}^{j}$ parametrizes a portion of the conditional projection $(\mathcal{P}_{t}^{j}f)(u):=\Pi^{j}f(u)+\int_{0}^{t}F_{t}^{j}(u,s)\,f(s)\,ds$, an orthogonal projection on $L^{2}([0,t];\mathbb{R}^{(n+1)d})$ (Remark A.4).
Lemma C.2 (Projection bounds).
For each $j$ and $t\geq 0$: (i) $\|\mathcal{P}_{t}^{j}\|_{\mathrm{op}}=1$; (ii) both $\mathcal{P}_{t}^{j}$ and $I-\mathcal{P}_{t}^{j}$ are $L^{2}$-contractions; (iii) $\|\mathcal{F}_{t}^{j}\|_{\mathrm{op}}\leq 2$ where $\mathcal{F}_{t}^{j}:=\mathcal{P}_{t}^{j}-\Pi^{j}$.
Proof.
Standard properties of orthogonal projections; (iii) is the triangle inequality. ∎
Lemma C.3 (Dissipation).
$\{\mathcal{P}_{t}^{j}\}_{t\geq 0}$ is non-decreasing in the positive semidefinite order, and $\|\widetilde{X}^{j}(t,\cdot)\|_{L^{2}([0,t])}\leq\|X(t,\cdot)\|_{L^{2}([0,t])}$ for all $t$.
Proof.
Equation (C.5) adds a positive semidefinite increment to $F_{t}^{j}$, so $\mathcal{P}_{t}^{j}$ is non-decreasing. Since $\widetilde{X}^{j}(t,\cdot)=(I-\mathcal{P}_{t}^{j})X(t,\cdot)$ by (C.2), Lemma C.2(ii) gives the bound. ∎
These bounds hold uniformly in the precision paths, the policy profile, and the horizon $T$. The quadratic right-hand side of (C.5) is self-limiting: as $\mathcal{P}_{t}^{j}$ grows toward $I$, the complementary projection contracts, driving $\widetilde{X}^{j}\to 0$ and decelerating $F^{j}$. This prevents finite-time blowup despite the quadratic nonlinearity and is the primitive noise-linear manifestation of Kalman variance collapse.
C.5 Well-posedness
Definition C.1.
$\mathbb{B}_{T}^{M}:=\bigl\{\mathbf{D}=\{(\bar{D}^{i},D^{i})\}_{i=1}^{n}:\;\bar{D}^{i}\in C([0,T];\mathbb{R}^{d}),\;D^{i}\in L^{\infty}(\Delta_{T}),\;\|\mathbf{D}\|\leq M\bigr\}$, where $\|\mathbf{D}\|:=\max_{i}(\|\bar{D}^{i}\|_{C}+\|D^{i}\|_{L^{\infty}})$.
Definition C.2 (Best-response operator).
$\mathcal{T}(\mathbf{D}):=\bigl\{(-(G^{DD,i})^{-1}\bar{H}^{X,i,\mathbf{D}},\;-(G^{DD,i})^{-1}H^{X,i,\mathbf{D}})\bigr\}_{i=1}^{n}$, where the adjoints solve the backward system of Theorem 3.2 at the forward environment induced by $\mathbf{D}$. Fixed points are Nash equilibria (Corollary 3.5, Remark 3.2).
The following estimates are proved by Grönwall arguments on the forward system (C.3)–(C.5) (using the a priori bounds of Lemmas C.2C.3 to control the quadratic nonlinearity in (C.5)) and by backward Grönwall on the linear adjoint system (B.11)–(B.14) (using Cauchy–Schwarz on the belief-adjoint coupling integrals $\int_{0}^{t}P^{k}\widetilde{X}^{k}\,\bar{H}^{k,i}\,dz$).
Proposition C.4 (Bounds and Lipschitz continuity).
Under Assumption C.1, for $\mathbf{D}\in\mathbb{B}_{T}^{M}$:
(i) The forward system has a unique global solution with $\|\bar{X}\|_{C}+\|X\|_{L^{\infty}}\leq K_{\mathrm{fwd}}$, $\|\widetilde{X}^{j}(t,\cdot)\|_{L^{2}}\leq K_{\mathrm{fwd}}\sqrt{T}$, $\|\mathcal{P}_{t}^{j}\|_{\mathrm{op}}=1$.
(ii) The backward system has a unique solution with $\|\bar{H}^{X,i}\|_{C}+\|H^{X,i}\|_{L^{\infty}}\leq K_{\mathrm{bwd}}$, and time derivatives bounded by $C(M)K_{\mathrm{bwd}}$.
(iii) Both maps $\mathbf{D}\mapsto(\bar{X},X,\widetilde{X}^{j})$ and $\mathbf{D}\mapsto(\bar{H}^{X,i},H^{X,i})$ are Lipschitz with constants $C_{\mathrm{fwd}}(T,M)$ and $C_{\mathrm{bwd}}(T,M)$ respectively, both bounded by $C(M)\,T\,e^{C(M)T}$.
Here $K_{\mathrm{fwd}},K_{\mathrm{bwd}}\leq C(M)e^{C(M)T}$.
Proof sketch.
For the forward system: $(\bar{X},X)$ solve linear equations with bounded coefficients once $\mathcal{D}^{k}$ is controlled. Since $\mathcal{D}^{k}$ involves $F^{k}$ through a linear integral, and $\|\mathcal{F}_{t}^{k}\|_{\mathrm{op}}\leq 2$ (Lemma C.2), the coupled system for $\phi(t):=\|X\|_{L^{\infty}(\Delta_{t})}$ closes as $\phi(t)\leq C(M)+C(M)\int_{0}^{t}\phi(r)\,dr$.
For the Lipschitz bound: the variation $\delta\mathcal{D}^{k}$ decomposes into a direct term bounded by $C(M)\|\delta\mathbf{D}\|$ and an indirect term through $\delta F^{k}$, which couples to $\delta\widetilde{X}^{k}$ and back to $\delta X$ via (C.2). A Grönwall argument on the aggregate $\Psi(t):=\|\delta X\|_{L^{\infty}(\Delta_{t})}+\max_{j}\sup_{r\leq t}\|\delta\widetilde{X}^{j}(r,\cdot)\|_{L^{2}}$ gives $\Psi(T)\leq C(M)Te^{C(M)T}\|\delta\mathbf{D}\|$; the factor of $T$ arises because the policy enters through integrals of length $\leq T$.
The backward estimates follow from linearity of the adjoint system, Cauchy–Schwarz on coupling terms, and backward Grönwall. ∎
Theorem C.5 (Short-horizon equilibrium).
Under Assumption C.1, there exists $T^{*}>0$ depending only on $(n,d,L,\lambda)$ such that for $\in(0,T^{*}], 𝒯\mathcal{T}$ is a contraction on $\mathbb{B}_{T}^{M_{0}}$ for a suitable $M_{0}$. In particular, there exists a unique Nash equilibrium in noise-state linear strategies with bounded kernels, Picard iterates converge geometrically, and the equilibrium is Nash over the full admissible $L^{2}$ class.
Proof.
By Proposition C.4(iii), $\mathrm{Lip}(\mathcal{T})\leq\lambda^{-1}C(M_{0})Te^{C(M_{0})T}$. Self-mapping ($\mathcal{T}:\mathbb{B}_{T}^{M_{0}}\to\mathbb{B}_{T}^{M_{0}}$) holds because at $M=0$ the belief-adjoint decouples ($D^{k}\equiv 0$ kills the forcing in (B.12)), giving finite $K_{\mathrm{bwd}}(0,T)$; the intermediate value theorem yields $M_{0}$. Since $M_{0}$ is bounded as $T\to 0$, choosing $T^{*}$ small enough makes $\mathrm{Lip}(\mathcal{T})\leq\tfrac{1}{2}$. Banach’s theorem gives existence, uniqueness, and geometric convergence. Corollary 3.5 extends to the full $L^{2}$ class. ∎
Theorem C.6 (Global existence).
Under Assumption C.1, for every $T>0$ there exists a Nash equilibrium over the full admissible $L^{2}$ class.
Proof sketch.
Apply the Schauder fixed-point theorem to $\mathcal{T}$ on $\mathbb{B}_{T}^{M_{0}}$, working in the topology $\tau:=(C([0,T])\times L^{2}(\Delta_{T}))^{n}$. Self-mapping holds by the same argument as above. $\tau$-closedness of $\mathbb{B}_{T}^{M_{0}}$: weak-* lower semicontinuity of the $L^{\infty}$ norm preserves the bound under $L^{2}$ limits. $\tau$-continuity of $\mathcal{T}$: on the bounded set $\mathbb{B}_{T}^{M_{0}}$, $L^{2}$-convergence of policy kernels propagates through the forward linear equations (via Cauchy–Schwarz) and the linear backward system. $\tau$-precompactness of the image: the uniform time-derivative bound on $\bar{H}^{X,i}$ (Proposition C.4(ii)) gives equicontinuity of $\bar{D}^{i,\mathrm{new}}$ in $C([0,T])$; the same bound gives pointwise-in-$u$ equicontinuity of $D^{i,\mathrm{new}}(t,u)$ in $t$, which with the uniform $L^{\infty}$ bound yields $L^{2}$-equicontinuity of $t\mapsto D^{i,\mathrm{new}}(t,\cdot)$ and hence sequential compactness in $L^{2}(\Delta_{T})$ by the vector-valued Arzelà–Ascoli theorem ([1, Theorem 1.3.2]). ∎
Remark C.1 (Uniqueness beyond the contraction regime).
Theorem C.6 gives existence but not uniqueness for $T>T^{*}$. The dissipation of Lemma C.3 suggests enough structure to preclude multiplicity, but a proof requires monotonicity estimates beyond the Grönwall bounds used here. In Section 4, Picard iteration converges to the same limit from multiple initial conditions for all parameter values tested.
Remark C.2 (Stochastic coefficients).
If $A(t,\omega)$, $P^{i}(t,\omega)$, or the cost matrices depend on a public factor process $\xi_{t}$ (e.g. a finite-state Markov chain representing regime switches or stochastic automation events), the kernel state augments to $(\mathcal{Z}_{t},\xi_{t})$. The projection bounds of Lemmas C.2C.3 are algebraic properties of orthogonal projections and hold pathwise, so the forward well-posedness extends to each sample path between jumps of $\xi$. Each jump triggers a regime transition in which filtering gains, strategies, and the information wedge shift discontinuously, while $\mathcal{Z}$ is continuous across jumps. The backward ODE is replaced by a backward stochastic differential equation in the adjoint kernels, driven by jumps of $\xi$; the deterministic Grönwall arguments extend to pathwise bounds conditional on the regime path.
Appendix D Kyle–Back derivations
This appendix derives the deviation propagation system, the first-order condition (5.3), and the closed backward ODE (5.4) stated in Section 5.
D.1 Drift kernels
The market maker’s drift kernel is $C^{0}(t,s):=\sigma_{0}^{-1}\mathcal{D}_{\Sigma,t}(s)$; for trader $k$:
$C^{k}(t,s):=\begin{pmatrix}\sigma_{0}^{-1}\mathcal{D}_{\Sigma,t}(s)\\[2.0pt] \sqrt{p_{k}}\,\sigma_{V}\,E^{V}\end{pmatrix}\!,\qquad\Pi_{\mathcal{E}}^{k}:=\Pi^{0}+\Pi^{k}.$ (D.1)
The unresolved kernels $\widetilde{C}^{0}$ and $\widetilde{C}^{k}$ are defined by the projection decomposition (A.12) with $C^{j}$ replacing $\sqrt{P^{j}}X$ and $\Pi_{\mathcal{E}}^{j}$ replacing $\Pi^{j}$.
D.2 Deviation propagation
A spike perturbation of trader $i$’s demand at time $t$ enters aggregate order flow, triggering a coupled forward response. For $s>t$, let $\delta D_{\Sigma,s}$ be the aggregate demand perturbation and $\delta w_{s}^{j}(u)$ the noise-state density perturbation for observer $j$. The deviation system (cf. Section B.2) is:
$\delta D_{\Sigma,s}$ $=\sum_{k\neq i}\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du,$ (D.2)
$\partial_{s}\delta w_{s}^{0}(u)$ $=\widetilde{C}^{0}(s,u)^{\top}\biggl(\sigma_{0}^{-1}\delta D_{\Sigma,s}-\int_{0}^{s}C^{0}(s,r)\,\delta w_{s}^{0}(r)\,dr\biggr),$ (D.3)
$\partial_{s}\delta w_{s}^{k}(u)$ $=\widetilde{C}^{k}(s,u)^{\top}\biggl(\begin{pmatrix}\sigma_{0}^{-1}\delta D_{\Sigma,s}\\ 0\end{pmatrix}-\int_{0}^{s}C^{k}(s,r)\,\delta w_{s}^{k}(r)\,dr\biggr),\quad k\neq i,$ (D.4)
$\delta P_{s}$ $=\sigma_{V}\int_{0}^{s}E^{V}\delta w_{s}^{0}(u)\,du,$ (D.5)
with $\delta w_{t}^{j}\equiv 0$, $\delta D_{\Sigma,t}=1$, $\delta I_{s}^{i}=0$. The demand perturbation enters each observer’s stacked observation through the order-flow row only. Equation (D.2) closes the loop: opponents’ noise-state shifts feed through their policy kernels back into aggregate demand. Since $E^{V}E^{0,\top}=0$, $\delta P_{t}=\lambda(t)$; when $n=1$, $\delta D_{\Sigma,s}=0$ for $s>t$ and $\delta P_{s}=\lambda(t)$ for all $s\geq t$.
D.3 First-order condition
The optimality condition from a unit demand impulse at $t$ is
$\mathbb{E}\!\left[V_{T}-P_{t}-\int_{t}^{T}D_{\tau}^{i}\,\delta P_{\tau}\,d\tau\;\Big|\;\mathcal{F}_{t}^{i}\right]=0.$ (D.6)
The mispricing is $\mathbb{E}[V_{T}-P_{t}\mid\mathcal{F}_{t}^{i}]=\sigma_{V}\int_{0}^{t}E^{V}\,d_{u}(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0})(u)$ (since $\mathcal{F}_{t}^{0}\subset\mathcal{F}_{t}^{i}$). The conditional future demand in noise-state linear form is $\mathbb{E}[D_{\tau}^{i}\mid\mathcal{F}_{t}^{i}]=\int_{0}^{t}D^{i}_{\tau}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$ for $\tau>t$ (the noise-state is a martingale, so only $u\leq t$ contributes). Both terms are linear in $\widehat{W}_{t}^{i}$, so matching noise-state densities gives (5.3). $D^{i}_{t}(u)$ does not appear directly: the integral over $[t,T]$ does not isolate $\tau=t$. To pin down $D^{i}_{t}(u)$, we differentiate both sides in $t$.
D.4 Leibniz differentiation and backward ODE
Impact adjoint.
Leibniz on $H^{i,\mathrm{imp}}$ gives
$\frac{d}{dt}H^{i,\mathrm{imp}}_{t}(u)=-\lambda(t)\,D^{i}_{t}(u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau,$ (D.7)
where the boundary term $\delta P_{t}\cdot D^{i}_{t}(u)=-\lambda(t)D^{i}_{t}(u)$ is instantaneous self-impact—this is where $\lambda(t)$ enters as the endogenous $G^{DD,i}$ and where $D^{i}_{t}(u)$ first appears explicitly.
Value adjoint.
In primitive-noise coordinates (Theorem 3.1(i)), $H^{i,\mathrm{val}}_{t}(u)$ has a $t$-independent term (the direct observation gap $\Pi_{\mathcal{E}}^{i}-\Pi^{0}$) and an integral $\int_{0}^{t}(F_{t}^{i}-F_{t}^{0})(v,u)\,dv$. Differentiating by Leibniz and the kernel evolution $\partial_{t}F_{t}^{j}(v,u)=\widetilde{C}^{j}(t,v)^{\top}\widetilde{C}^{j}(t,u)$:
$\frac{d}{dt}H^{i,\mathrm{val}}_{t}(u)=\sigma_{V}E^{V}\biggl[\bigl(F_{t}^{i}-F_{t}^{0}\bigr)(t,u)+\int_{0}^{t}\bigl(\widetilde{C}^{i}(t,v)^{\top}\widetilde{C}^{i}(t,u)-\widetilde{C}^{0}(t,v)^{\top}\widetilde{C}^{0}(t,u)\bigr)\,dv\biggr].$ (D.8)
Closed backward ODE.
Equating $\frac{d}{dt}H^{i,\mathrm{val}}=\frac{d}{dt}H^{i,\mathrm{imp}}$ and solving for $D^{i}_{t}(u)$ yields (5.4). The derivative $\partial_{t}\delta P_{\tau}$ is obtained by differentiating (D.2)–(D.5) in the impulse time, following the same logic as Section B.2 with $\sqrt{P^{k}}X\to C^{k}$, $\widetilde{X}^{k}\to\widetilde{C}^{k}$. Convexity (from $\lambda(t)>0$) ensures sufficiency.